AUCTEX

A sophisticated TEX environment for Emacs Version 11.84, 2007-01-12

Kresten Krab Thorup Per Abrahamsen David Kastrup and others

This manual is for AUCTEX (version 11.84 from 2007-01-12), a sophisticated TeX environment for Emacs. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1992, 1993, 1994, 1995, 2001, 2002, 2004, 2005, 2006 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.2 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover Texts and no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License.”

Table of Contents

Executive Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Copying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

1 Introduction to AUCTeX. . . . . . . . . . . . . . . . . . . 3

1.1 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.2 Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Availability . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.4 Contacts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 Installing AUCTeX . . . . . . . . . . . . . . . . . . . . . . . . 5

2.1 Prerequisites . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.2 Configure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.3 Build/install . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.4 Loading the package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.5 Providing AUCTeX as a package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.6 Installation for non-privileged users . . . . . . . . . . . . . . . . . . . . . . . . . . 10 2.7 Installation under MS Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.8 Customizing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

3 Quick Start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

3.1 Functions for editing TeX files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.1.1 Making your TeX code more readable . . . . . . . . . . . . . . . . . . . 17 3.1.2 Entering sectioning commands . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.1.3 Inserting environments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.1.4 Inserting macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.1.5 Changing the font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.1.6 Other useful features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

3.2 Creating and viewing output, debugging . . . . . . . . . . . . . . . . . . . . . 19 3.2.1 One Command for LaTeX, helpers, viewers, and printing. . 19 3.2.2 Choosing an output format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.2.3 Debugging LaTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.2.4 Running LaTeX on parts of your document . . . . . . . . . . . . . . 20

4 Inserting Frequently Used Commands. . . . . . 21

4.1 Insertion of Quotes, Dollars, and Braces. . . . . . . . . . . . . . . . . . . . . . 21 4.2 Inserting Font Specifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.3 Inserting chapters, sections, etc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.4 Inserting Environment Templates . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

4.4.1 Equations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 4.4.2 Floats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.4.3 Itemize-like. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.4.4 Tabular-like . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.4.5 Customizing environments. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

5 Advanced Editing Features . . . . . . . . . . . . . . . . 28

5.1 Entering Mathematics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 5.2 Completion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 5.3 Commenting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 5.4 Indenting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 5.5 Filling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

6 Controlling Screen Display . . . . . . . . . . . . . . . . 35

6.1 Font Locking. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 6.2 Folding Macros and Environments . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 6.3 Outlining the Document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

7 Starting Processors, Viewers and Other

Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

7.1 Executing Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 7.2 Viewing the formatted output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

7.2.1 Starting viewers and customizing their invocation . . . . . . . . 47 7.2.2 Forward and inverse search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

7.3 Catching the errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 7.4 Checking for problems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 7.5 Controlling the output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 7.6 Cleaning intermediate and output files . . . . . . . . . . . . . . . . . . . . . . . 49 7.7 Documentation about macros and packages . . . . . . . . . . . . . . . . . . 50

8 Multifile Documents . . . . . . . . . . . . . . . . . . . . . . 51

9 Automatic Parsing of TeX Files . . . . . . . . . . . 53

10 Language Support . . . . . . . . . . . . . . . . . . . . . . . 55

10.1 Using AUCTeX with European Languages . . . . . . . . . . . . . . . . . . 55 10.1.1 Typing and Displaying Non-ASCII Characters . . . . . . . . . . 55 10.1.2 Style Files for Different Languages . . . . . . . . . . . . . . . . . . . . . 55

10.2 Using AUCTeX with Japanese TeX . . . . . . . . . . . . . . . . . . . . . . . . . 57

11 Automatic Customization . . . . . . . . . . . . . . . . 59

11.1 Automatic Customization for the Site. . . . . . . . . . . . . . . . . . . . . . . 59 11.2 Automatic Customization for a User . . . . . . . . . . . . . . . . . . . . . . . . 60 11.3 Automatic Customization for a Directory . . . . . . . . . . . . . . . . . . . 60

12 Writing Your own Style Support. . . . . . . . . . 62

12.1 A Simple Style File. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 12.2 Adding Support for Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 12.3 Adding Support for Environments . . . . . . . . . . . . . . . . . . . . . . . . . . 65 12.4 Adding Other Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 12.5 Automatic Extraction of New Things . . . . . . . . . . . . . . . . . . . . . . . 66

Appendix A Copying this Manual . . . . . . . . . . . 69

A.1 GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . . 69

Appendix B Changes and New Features . . . . . 76

Appendix C Future Development . . . . . . . . . . . 84

C.1 Mid-term Goals. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 C.2 Wishlist. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 C.3 Bugs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

Appendix D Frequently Asked Questions …. 88

Key Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

Function Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

Variable Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

Executive Summary

AUCTEX is an integrated environment for editing LaTEX, ConTEXt, docTEX, Texinfo, and TEX files.

Although AUCTEX contains a large number of features, there are no reasons to despair. You can continue to write TEX and LaTEX documents the way you are used to, and only start using the multiple features in small steps. AUCTEX is not monolithic, each feature described in this manual is useful by itself, but together they provide an environment where you will make very few LaTEX errors, and makes it easy to find the errors that may slip through anyway.

It is a good idea to make a printout of AUCTEX’s reference card ‘tex-ref.tex’ or one of its typeset versions.

If you want to make AUCTEX aware of style files and multi-file documents right away, insert the following in your ‘.emacs’ file.

     (setq TeX-auto-save t)
     (setq TeX-parse-self t)
     (setq-default TeX-master nil)

Another thing you should enable is RefTEX, a comprehensive solution for managing cross references, bibliographies, indices, document navigation and a few other things. (see section “Installation” in The RefTEX manual)

For detailed information about the preview-latex subsystem of AUCTEX, see section “Introduction” in The preview-latex Manual.

There is a mailing list for general discussion about AUCTEX: write a mail with “subscribe” in the subject to auctex-request@gnu.org to join it. Send contributions to

auctex@gnu.org.

Bug reports should go to bug-auctex@gnu.org, suggestions for new features, and pleas for help should go to either auctex-devel@gnu.org (the AUCTEX developers), or to auctex@gnu.org if they might have general interest. Please use the command M-x TeX-submit-bug-report RET to report bugs if possible. You can subscribe to a low-volume announcement list by sending “subscribe” in the subject of a mail to

info-auctex-request@gnu.org.

Copying

AUCTEX primarily consists of Lisp files for Emacs (and XEmacs), but there are also installation scripts and files and TEX support files. All of those are free; this means that everyone is free to use them and free to redistribute them on a free basis. The files of AUCTEX are not in the public domain; they are copyrighted and there are restrictions on their distribution, but these restrictions are designed to permit everything that a good cooperating citizen would want to do. What is not allowed is to try to prevent others from further sharing any version of these programs that they might get from you.

Specifically, we want to make sure that you have the right to give away copies of the files that constitute AUCTEX, that you receive source code or else can get it if you want it, that you can change these files or use pieces of them in new free programs, and that you know you can do these things.

To make sure that everyone has such rights, we have to forbid you to deprive anyone else of these rights. For example, if you distribute copies of parts of AUCTEX, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must tell them their rights.

Also, for our own protection, we must make certain that everyone finds out that there is no warranty for AUCTEX. If any parts are modified by someone else and passed on, we want their recipients to know that what they have is not what we distributed, so that any problems introduced by others will not reflect on our reputation.

The precise conditions of the licenses for the files currently being distributed as part of AUCTEX are found in the General Public Licenses that accompany them. This manual specifically is covered by the GNU Free Documentation License (see Appendix A [Copying this Manual], page 69).

1 Introduction to AUCTEX

This section of the AUCTEX manual gives a brief overview of what AUCTEX is. It is not an attempt to document AUCTEX. Real documentation for AUCTEX is available in the rest of the manual.

1.1 Installation

Read the section Chapter 2 [Installation], page 5, or Section 2.7 [Installation under MS Windows], page 11, respectively for comprehensive information about how to install AUCTEX.

The installation routine tries to make the modes provided by AUCTEX the default for all supported file types. If this does not happen in your case, add

     (load "auctex.el" nil t t)

to your init file and consult the section Section 2.4 [Loading the package], page 8.

If you want to change the modes for which it is operative instead of the default, use

     M-x customize-variable RET TeX-modes RET

If you want to remove a preinstalled AUCTEX completely before any of its modes have been used,

     (unload-feature 'tex-site)

should accomplish that.

If you are considering upgrading AUCTEX, the recent changes are described in Appendix B [Changes], page 76.

1.2 Features

AUCTEX is a comprehensive customizable integrated environment for writing input files for TEX/LaTEX/ConTEXt/Texinfo using Emacs or XEmacs.

It lets you process your source files by running TEX and related tools (such as output filters, post processors for generating indices and bibliographies, and viewers) from inside Emacs. AUCTEX lets you browse through the errors TEX reported, while it moves the cursor directly to the reported error, and displays some documentation for that particular error. This will even work when the document is spread over several files.

One component of AUCTEX that LaTEX users will find attractive is preview-latex, a combination of folding and in-source previewing that provides true “What You See Is What You Get” experience in your sourcebuffer, while letting you retain full control. preview-latex comes with its own manual, see section “preview-latex” in The preview-latex Manual.

AUCTEX automatically indents your ‘LaTEX-source’, not only as you write it — you can also let it indent and format an entire document. It has a special outline feature, which can greatly help you ‘getting an overview’ of a document.

Apart from these special features, AUCTEX provides a large range of handy Emacs macros, which in several different ways can help you write your documents fast and painlessly.

All features of AUCTEX are documented using the GNU Emacs online documentation system. That is, documentation for any command is just a key click away!

AUCTEX is written entirely in Emacs-Lisp, and hence you can easily add new features for your own needs. It has become recently a GNU project. AUCTEX is distributed under the ‘GNU General Public License Version 2’.

1.3 Availability

The most recent version is always available at

                                         http://ftp.gnu.org/pub/gnu/auctex/

WWW users may want to check out the AUCTEX page at

                                        http://www.gnu.org/software/auctex/

1.4 Contacts

Various mailing lists exist.

Send a mail with the subject “subscribe” to auctex-request@gnu.org in order to join the general discussion list for AUCTEX. Articles should be sent to auctex@gnu.org. In a similar way, you can subscribe to the info-auctex@gnu.org list for just getting important announcements about AUCTEX. The list bug-auctex@gnu.org is for bug reports which you should usually file with the M-x TeX-submit-bug-report RET command.

If you want to address the developers of AUCTEX themselves with technical issues, they can be found on the discussion list auctex-devel@gnu.org.

2 Installing AUCTEX

Installing AUCTEX should be simple: merely ./configure, make, and make install for a standard site-wide installation (most other installations can be done by specifying a

‘--prefix=...’ option).

On many systems, this will already activate the package, making its modes the default instead of the built-in modes of Emacs. If this is not the case, consult Section 2.4 [Loading the package], page 8. Please read through this document fully before installing anything. The installation procedure has changed as compared to earlier versions. Users of MS Windows are asked to consult See Section 2.7 [Installation under MS Windows], page 11.

2.1 Prerequisites

• A recent version of Emacs, alternatively XEmacs Emacs 20 is no longer supported, and neither is XEmacs with a version of xemacsbase older than 1.84 (released in sumo from 02/02/2004). Using preview-latex requires a version of Emacs compiled with image support. This means that Emacs 21 will work only in the version for X11: for Windows and Mac OS X, you need to use Emacs 22 (which is not yet released) or a developer version. Since the developer version is quite stable by now and features five more years of development and bugfixes, we recommend its use even for X11-based platforms. You can get it here:

Windows Precompiled ver-

          sions are available from ftp://alpha.gnu.org/gnu/auctex/ (including
          AUCTEX) and http://ourcomments.org/Emacs/EmacsW32.html.

Mac OS X A precompiled version including an installer as well as preinstalled versions of AUCTEX and preview-latex is available from

          http://homepage.mac.com/zenitani/emacs-e.html.

Debian GNU/Linux

          Debian provides ‘emacs-snapshot’ and ‘emacs-snapshot-gtk’ packages in

its ‘unstable’ distribution.

Fedora GNU/Linux

          http://people.redhat.com/coldwell/bugs/emacs/176171/

Self-compiled Compiling Emacs yourself requires a C compiler and a number of tools and development libraries. Details are beyond the scope of this manual. Instructions for checking out the source code can be found at

          http://savannah.gnu.org/cvs/?group=emacs.

If you really need to use Emacs 21 on platforms where this implies missing image support, you should disable the installation of preview-latex (see below). While XEmacs (version 21.4.15, 21.4.17 or later) is supported, doing this in a satisfactory manner has proven to be difficult. This is mostly due to technical shortcomings and differing API’s which are hard to come by. If AUCTEX is your main application for XEmacs, you are likely to get better results and support by switching to Emacs. Of course, you can improve support for your favorite editor by giving feedback in case you encounter bugs.

• A working TEX installation Well, AUCTEX would be pointless without that. Processing documentation requires TEX, LaTEX and Texinfo during installation. preview-latex requires Dvips for its operation in DVI mode. The default configuration of AUCTEX is tailored for teTEX-based distributions, but can be adapted easily.

• A recent Ghostscript This is needed for operation of preview-latex in both DVI and PDF mode. Most versions of Ghostscript nowadays in use should work fine (version 7.0 and newer). If you encounter problems, check section “Problems with Ghostscript” in the preview-latex manual.

• The texinfo package Strictly speaking, you can get away without it if you are building from the distribution tarball, have not modified any files and don’t need a printed version of the manual: the pregenerated info file is included in the tarball. At least version 4.0 is required.

For some known issues with various software, see section “Known problems” in the preview-latex manual.

2.2 Configure

The first step is to configure the source code, telling it where various files will be. To do so, run

./configure options

(Note: if you have fetched AUCTEX from CVS rather than a regular release, you will have to first follow the instructions in ‘README.CVS’).

On many machines, you will not need to specify any options, but if configure cannot determine something on its own, you’ll need to help it out with one of these options:

--prefix=`/usr/local'

All automatic placements for package components will be chosen from sensible existing hierarchies below this: directories like ‘man’, ‘share’ and ‘bin’ are supposed to be directly below prefix. Only if no workable placement can be found there, in some cases an alternative search will be made in a prefix deduced from a suitable binary. ‘/usr/local’ is the default prefix, intended to be suitable for a site-wide installation. If you are packaging this as an operating system component for distribution, the setting ‘/usr’ will probably be the right choice. If you are planning to install the package as a single non-priviledged user, you will typically set prefix to your home directory.

--with-emacs[=/path/to/emacs]

If you are using a pretest which isn’t in your $PATH, or configure is not finding the right Emacs executable, you can specify it with this option.

--with-xemacs[=/path/to/xemacs]

Configure for generation under XEmacs (Emacs is the default). Again, the name of the right XEmacs executable can be specified, complete with path if necessary.

--with-packagedir=/dir

This XEmacs-only option configures the directory for XEmacs packages. A typical user-local setting would be ‘~/.xemacs/xemacs-packages’. If this directory exists and is below prefix, it should be detected automatically. This will install and activate the package.

--without-packagedir

This XEmacs-only option switches the detection of a package directory and corresponding installation off. Consequently, the Emacs installation scheme will be used. This might be appropriate if you are using a different package system/installer than the XEmacs one and want to avoid conflicts. The Emacs installation scheme has the following options:

--with-lispdir=/dir

This Emacs-only option specifies the location of the ‘site-lisp’ directory within ‘load-path’ under which the files will get installed (the bulk will get installed in a subdirectory). ‘./configure’ should figure this out by itself.

--with-auctexstartfile=`auctex.el'
--with-previewstartfile=`preview-latex.el'

This is the name of the respective startup files. If lispdir contains a subdirectory ‘site-start.d’, the start files are placed there, and ‘site-start.el’ should load them automatically. Please be aware that you must not move the start files after installation since other files are found relative to them.

--with-packagelispdir=`auctex'

This is the directory where the bulk of the package gets located. The startfile adds this into load-path.

--with-auto-dir=/dir

You can use this option to specify the directory containing automatically generated information. It is not necessary for most TEX installs, but may be used if you don’t like the directory that configure is suggesting.

--help This is not an option specific to AUCTEX. A number of standard options to configure exist, and we do not have the room to describe them here; a short description of each is available, using --help. If you use ‘--help=recursive’, then also preview-latex-specific options will get listed.

--disable-preview

This disables configuration and installation of preview-latex. This option is not actually recommended. If your Emacs does not support images, you should really upgrade to a newer version. Distributors should, if possible, refrain from distributing AUCTEX and preview-latex separately in order to avoid confusion and upgrade hassles if users install partial packages on their own.

--with-texmf-dir=/dir
--without-texmf-dir

This option is used for specifying a TDS-compliant directory hierarchy. Using -- with-texmf-dir=/dir you can specify where the TEX TDS directory hierarchy resides, and the TEX files will get installed in ‘/dir/tex/latex/preview/’.

If you use the --without-texmf-dir option, the TEX-related files will be kept in the Emacs Lisp tree, and at runtime the TEXINPUTS environment variable will be made to point there. You can install those files into your own TEX tree

          at some later time with M-x preview-install-styles RET.
--with-tex-dir=/dir

If you want to specify an exact directory for the preview TEX files, use --with-tex-dir=/dir. In this case, the files will be placed in ‘/dir’, and you’ll also need the following option:

--with-doc-dir=/dir

This option may be used to specify where the TEX documentation goes. It is to be used when you are using --with-tex-dir=/dir, but is normally not necessary otherwise.

2.3 Build/install

Once configure has been run, simply enter

     make

at the prompt to byte-compile the lisp files, extract the TEX files and build the documentation files. To install the files into the locations chosen earlier, type

     make install

You may need special privileges to install, e.g., if you are installing into system directories.

2.4 Loading the package

You can detect the successful activation of AUCTEX and preview-latex in the menus after loading a LaTEX file like ‘preview/circ.tex’: AUCTEX then gives you a ‘Command’ menu, and preview-latex gives you a ‘Preview’ menu.

For XEmacs, if the installation occured into a valid package directory (which is the default), then this should work out of the box.

With Emacs (or if you explicitly disabled use of the package system), the startup files

‘auctex.el’ and ‘preview-latex.el’ may already be in a directory of the ‘site-start.d/’

variety if your Emacs installation provides it. In that case they should be automatically loaded on startup and nothing else needs to be done. If not, they should at least have been placed somewhere in your load-path. You can then load them by placing the lines

     (load "auctex.el" nil t t)
     (load "preview-latex.el" nil t t)

into your ‘~/.emacs’ file.

If you explicitly used --with-lispdir, you may need to add the specified directory into Emacs’ load-path variable by adding something like

     (add-to-list 'load-path "~/elisp")

before the above lines into your Emacs startup file.

For site-wide activation in GNU Emacs, see See Section 2.5 [Advice for package providers], page 9.

That is all. There are other ways of achieving the equivalent thing, but we don’t mention them here any more since they are not better, and people got confused into trying everything at once.

2.5 Providing AUCTEX as a package

As a package provider, you should make sure that your users will be served best according to their intentions, and keep in mind that a system might be used by more than one user, with different preferences.

There are people that prefer the built-in Emacs modes for editing TEX files, in particular plain TEX users. There are various ways to tell AUCTEX even after auto-activation that it should not get used, and they are described in Chapter 1 [Introduction to AUCTEX], page 3.

So if you have users that don’t want to use the preinstalled AUCTEX, they can easily get rid of it. Activating AUCTEX by default is therefore a good choice.

If the installation procedure did not achieve this already by placing ‘auctex.el’ and ‘preview-latex.el’ into a possibly existing ‘site-start.d’ directory, you can do this by placing

     (load "auctex.el" nil t t)
     (load "preview-latex.el" nil t t)

in the system-wide ‘site-start.el’.

If your package is intended as an XEmacs package or to accompany a precompiled version of Emacs, you might not know which TEX system will be available when preview-latex gets used. In this case you should build using the --without-texmf-dir option described previously. This can also be convenient for systems that are intended to support more than a single TeX distribution. Since more often than not TEX packages for operating system distributions are either much more outdated or much less complete than separately provided systems like TEX Live, this method may be generally preferable when providing packages.

The following package structure would be adequate for a typical fully supported Unix-like installation:

‘preview-tetex’

Style files and documentation for ‘preview.sty’, placed into a TEX tree where it is accessible from the teTEX executables usually delivered with a system. If there are other commonly used TEX system packages, it might be appropriate to provide separate packages for those.

‘auctex-emacs-tetex’

This package will require the installation of ‘preview-tetex’ and will record in ‘TeX-macro-global’ where to find the TEX tree. It is also a good idea to run

               emacs -batch -f TeX-auto-generate-global

when either AUCTEX or teTEX get installed or upgraded. If your users might want to work with a different TEX distribution (nowadays pretty common), instead consider the following:

‘auctex-emacs’

This package will be compiled with ‘--without-texmf-dir’ and will consequently contain the ‘preview’ style files in its private directory. It will probably not be possible to initialize ‘TeX-macro-global’ to a sensible value, so running ‘TeX-auto-generate-global’ does not appear useful. This package would neither conflict with nor provide ‘preview-tetex’.

‘auctex-xemacs-tetex’
‘auctex-xemacs’

Those are the obvious XEmacs equivalents. For XEmacs, there is the additional problem that the XEmacs sumo package tree already possibly provides its own version of AUCTEX, and the user might even have used the XEmacs package manager to updating this package, or even installing a private AUCTEX version. So you should make sure that such a package will not conflict with existing XEmacs packages and will be at an appropriate place in the load order (after site-wide and user-specific locations, but before a distribution-specific sumo package tree). Using the --without-packagedir option might be one idea to avoid conflicts. Another might be to refrain from providing an XEmacs package and just rely on the user or system administrator to instead use the XEmacs package system.

2.6 Installation for non-privileged users

Often people without system administration privileges want to install software for their private use. In that case you need to pass more options to the configure script. For XEmacs users, this is fairly easy, because the XEmacs package system has been designed to make this sort of thing practical: but GNU Emacs users (and XEmacs users for whom the package system is for some reason misbehaving) may need to do a little more work.

The main expedient is using the ‘--prefix’ option to the ‘configure’ script, and let it point to the personal home directory. In that way, resulting binaries will be installed under the ‘bin’ subdirectory of your home directory, manual pages under ‘man’ and so on. It is reasonably easy to maintain a bunch of personal software, since the prefix argument is supported by most ‘configure’ scripts.

You’ll have to add something like ‘/home/myself/share/emacs/site-lisp’ to your load-path variable, if it isn’t there already.

XEmacs users can achieve the same end by pointing configure at an appropriate package

directory (normally ‘--with-packagedir=~/.xemacs/xemacs-packages’ will serve). The

package directory stands a good chance at being detected automatically as long as it is in a subtree of the specified prefix.

Now here is another thing to ponder: perhaps you want to make it easy for other users to share parts of your personal Emacs configuration. In general, you can do this by writing ‘~myself/’ anywhere where you specify paths to something installed in your personal subdirectories, not merely ‘~/’, since the latter, when used by other users, will point to non-existent files.

For yourself, it will do to manipulate environment variables in your ‘.profile’ resp. ‘.login’ files. But if people will be copying just Elisp files, their copies will not work. While it would in general be preferable if the added components where available from a shell level, too (like when you call the standalone info reader, or try using ‘preview.sty’ for functionality besides of Emacs previews), it will be a big help already if things work from inside of Emacs.

Here is how to do the various parts:

Making the Elisp available

In GNU Emacs, it should be sufficient if people just do

     (load "~myself/share/emacs/site-lisp/auctex.el" nil t t)
     (load "~myself/share/emacs/site-lisp/preview-latex.el" nil t t)

where the path points to your personal installation. The rest of the package should be found relative from there without further ado.

In XEmacs, you should ask the other users to add symbolic links in the subdirectories ‘lisp’, ‘info’ and ‘etc’ of their ‘~/.xemacs/xemacs-packages/’ directory. (Alas, there is presently no easy programmatic way to do this, except to have a script do the symlinking for them.)

Making the Info files available

For making the info files accessible from within Elisp, something like the following might be convenient to add into your or other people’s startup files:

     (eval-after-load 'info
        '(add-to-list 'Info-directory-list "~myself/info"))

In XEmacs, as long as XEmacs can see the package, there should be no need to do anything at all; the info files should be immediately visible. However, you might want to set INFOPATH anyway, for the sake of standalone readers outside of XEmacs. (The info files

in XEmacs are normally in ‘~/.xemacs/xemacs-packages/info’.)

Making the LaTEX style available

If you want others to be able to share your installation, you should configure it using ‘--without-texmf-dir’, in which case things should work as well for them as for you.

2.7 Installation under MS Windows

In a Nutshell

The following are brief installation instructions for the impatient. In case you don’t understand some of this, run into trouble of some sort, or need more elaborate information, refer to the detailed instructions further below.

1. Install the prerequisites, i.e. GNU Emacs or XEmacs, MSYS or Cygwin, a TEX system, and Ghostscript.

2. Open the MSYS shell or a Cygwin shell and change to the directory containing the unzipped file contents.

3. Configure AUCTEX:

For GNU Emacs: Many people like to install AUCTEX into the pseudo file system hierarchy set up by the Emacs installation. Assuming Emacs is installed in ‘C:/Program Files/Emacs’ and the directory for local additions of your TEX system, e.g. MiKTEX, is ‘C:/localtexmf’, you can do this by typing the following statement at the shell prompt:

         ./configure --prefix='C:/Program Files/Emacs' \
           --with-texmf-dir='C:/localtexmf'

For XEmacs: You can install AUCTEX as an XEmacs package. Assuming XEmacs is installed in ‘C:/Program Files/XEmacs’ and the directory for local additions of your TEX system, e.g. MiKTEX, is ‘C:/localtexmf’, you can do this by typing the following command at the shell prompt:

         ./configure --with-xemacs='C:/Program Files/XEmacs/bin/xemacs' \
           --with-texmf-dir='C:/localtexmf'

The commands above are examples for common usage. More on configuration options can be found in the detailed installation instructions below. If the configuration script failed to find all required programs, make sure that these programs are in your system path and add directories containing the programs to the PATH environment variable if necessary. Here is how to do that in W2000/XP:

1. On the desktop, right click “My Computer” and select properties.

2. Click on “Advanced” in the “System Properties” window.

3. Select “Environment Variables”.

4. Select “path” in “System Variables” and click “edit”. Move to the front in the line (this might require scrolling) and add the missing path including drive letter, ended with a semicolon.

4. If there were no further error messages, type

         make

In case there were, please refer to the detailed description below.

5. Finish the installation by typing

         make install

Detailed Installation Instructions

Installation of AUCTEX under Windows is in itself not more complicated than on other platforms. However, meeting the prerequisites might require more work than on some other platforms, and feel less natural.

If you are experiencing any problems, even if you think they are of your own making, be sure to report them to auctex-devel@gnu.org so that we can explain things better in future.

Windows is a problematic platform for installation scripts. The main problem is that the installation procedure requires consistent file names in order to find its way in the directory hierarchy, and Windows path names are a mess.

The installation procedure tries finding stuff in system search paths and in Emacs paths. For that to succeed, you have to use the same syntax and spelling and case of paths everywhere: in your system search paths, in Emacs’ load-path variable, as argument to the scripts. If your path names contain spaces or other ‘shell-unfriendly’ characters, most notably backslashes for directory separators, place the whole path in ‘"double quote marks"’ whenever you specify it on a command line.

Avoid ‘helpful’ magic file names like ‘/cygdrive/c’ and ‘C:\PROGRA~1\’ like the plague. It is quite unlikely that the scripts will be able to identify the actual file names involved. Use the full paths, making use of normal Windows drive letters like ‘ 'C:/Program Files/Emacs' ’ where required, and using the same combination of upper- and lowercase letters as in the actual files. File names containing shell-special characters like spaces or backslashes (if you prefer that syntax) need to get properly quoted to the shell: the above example used single quotes for that.

Ok, now here are the steps to perform:

1. You need to unpack the AUCTEX distribution (which you seemingly have done since you are reading this). It must be unpacked in a separate installation directory outside of your Emacs file hierarchy: the installation will later copy all necessary files to their final destination, and you can ultimately remove the directory where you unpacked the files.

Line endings are a problem under Windows. The distribution contains only text files, and theoretically most of the involved tools should get along with that. However, the files are processed by various utilities, and it is conceivable that not all of them will use the same line ending conventions. If you encounter problems, it might help if you try unpacking (or checking out) the files in binary mode, if your tools allow that.

If you don’t have a suitable unpacking tool, skip to the next step: this should provide you with a working ‘unzip’ command.

2. The installation of AUCTEX will require the MSYS tool

set from http://www.mingw.org or the Cygwin tool set from http://cygwin.com.

The latter is slower and larger (the download size of the base system is about 15 MB) but comes with a package manager that allows for updating the tool set and installing additional packages like, for example, the spell checker aspell.

If Cygwin specific paths like ‘/cygdrive/c’ crop up in the course of the installation, using a non-Cygwin Emacs could conceivably cause trouble. Using Cygwin either for everything or nothing might save headaches, if things don’t work out.

3. Install a current version of XEmacs from http://www.xemacs.org or try getting and

compiling  a  developer

version of Emacs 22 from Savannah (http://savannah.gnu.org/cvs/?group=emacs).

Compiling Emacs is outside of the scope of this manual. Precompiled versions are available from ftp://alpha.gnu.org/gnu/auctex/ (including AUCTEX) and

http://ourcomments.org/Emacs/EmacsW32.html.

If you don’t want to use a developer version and Emacs 22 has not yet been released, it is also possible to use an Emacs 21 binary from

http://ftp.gnu.org/pub/gnu/windows/emacs/, but then you should disable the in-

stallation of preview-latex (it will not work). Since the developer version has seen quite a few improvements relevant also for other features of AUCTEX, we really recommend you give it a try.

4. You need a working TEX installation. One popular installation under Windows is MiKTEX (http://www.miktex.org). Another much more extensive system is TEX Live (http://www.tug.org/texlive) which is rather close to its Unix cousins.

5. A working copy of Ghostscript (http://www.cs.wisc.edu/~ghost) is required for preview-latex operation. Examining the output from

     gswin32c -h

on a Windows command line should tell you whether your Ghostscript supports the png16m device needed for PNG support. MiKTeX apparently comes with its own Ghostscript called ‘mgs.exe’.

6. Perl (http://www.perl.org) is needed for rebuilding the documentation if you are working with a copy from CVS or have touched documentation source files in the preview-latex part. If the line endings of the file ‘preview/latex/preview.dtx’ don’t correspond with what Perl calls \n when reading text files, you’ll run into trouble.

7. Now the fun stuff starts. If you have not yet done so, unpack the AUCTEX distribution into a separate directory after rereading the instructions for unpacking above.

8. Ready for takeoff. Start some shell (typically bash) capable of running configure, change into the installation directory and call ./configure with appropriate options. Typical options you’ll want to specify will be

--prefix=drive:/path/to/emacs-hierarchy

which tells ‘configure’ where to perform the installation. It may also make ‘configure’ find Emacs or XEmacs automatically; if this doesn’t happen, try one of ‘--with-emacs’ or ‘--with-xemacs’ as described below. All automatic detection of files and directories restricts itself to directories below the prefix or in the same hierarchy as the program accessing the files. Usually, directories like ‘man’, ‘share’ and ‘bin’ will be situated right under prefix. This option also affects the defaults for placing the Texinfo documentation files and automatically generated style hooks. If you have a central directory hierarchy (not untypical with Cygwin) for such stuff, you might want to specify its root here. You stand a good chance that this will be the only option you need to supply, as long as your TEX-related executables are in your system path, which they better be for AUCTEX’s operation, anyway.

--with-emacs

if you are installing for a version of Emacs. You can use

‘--with-emacs=drive:/path/to/emacs’ to specify the name of the in-

stalled Emacs executable, complete with its path if necessary (if Emacs is not within a directory specified in your PATH environment setting).

--with-xemacs

if you are installing for a version of XEmacs. Again, you can use

‘--with-xemacs=drive:/path/to/xemacs’ to specify the name of the in-

stalled XEmacs executable complete with its path if necessary. It may also be necessary to specify this option if a copy of Emacs is found in your PATH environment setting, but you still would like to install a copy of AUCTEX for XEmacs.

--with-packagedir=drive:/dir

is an XEmacs-only option giving the location of the package directory. This will install and activate the package. Emacs uses a different installation scheme:

--with-lispdir=drive:/path/to/site-lisp

This Emacs-only option tells a place in load-path below which the files are situated. The startup files ‘auctex.el’ and ‘preview-latex.el’ will get installed here unless a subdirectory ‘site-start.d’ exists which will then be used instead. The other files from AUCTEX will be installed in a subdirectory called ‘auctex’. If you think that you need a different setup, please refer to the full installation instructions in Section 2.2 [Configure], page 6.

--with-auto-dir=drive:/dir

Directory containing automatically generated information. You should not normally need to set this, as ‘--prefix’ should take care of this.

--disable-preview

Use this option if your Emacs version is unable to support image display. This will be the case if you are using a native variant of Emacs 21.

--with-texmf-dir=drive:/dir

This will specify the directory where your TEX installation sits. If your TEX installation does not conform to the TDS (TEX directory standard), you may need to specify more options to get everything in place:

For more information about any of the above and additional options, see Section 2.2 [Configure], page 6. Calling ‘./configure --help=recursive’ will tell about other options, but those are almost never required. Some executables might not be found in your path. That is not a good idea, but you can get around by specifying environment variables to ‘configure’:

     GS="drive:/path/to/gswin32c.exe" ./configure ...

should work for this purpose. ‘gswin32c.exe’ is the usual name for the required command line executable under Windows; in contrast, ‘gswin32.exe’ is likely to fail. As an alternative to specifying variables for the ‘configure’ call you can add directories containing the required executables to the PATH variable of your Windows system. This is especially a good idea if Emacs has trouble finding the respective programs later during normal operation.

9. Run make in the installation directory.

10. Run make install in the installation directory.

11. With XEmacs, AUCTEX and preview-latex should now be active by default. With Emacs, activation depends on a working ‘site-start.d’ directory or similar setup, since then the startup files ‘auctex.el’ and ‘preview-latex.el’ will have been placed there. If this has not been done, you should be able to load the startup files manually with

     (load "auctex.el" nil t t)
     (load "preview-latex.el" nil t t)

in either a site-wide ‘site-start.el’ or your personal startup file (usually accessible as ‘~/.emacs’ from within Emacs and ‘~/.xemacs/init.el’ from within XEmacs).

The default configuration of AUCTEX is probably not the best fit for Windows systems. You might want to additionally use

         (require 'tex-mik)

or

         (require 'tex-fptex)

in order to get more appropriate values for MiKTEX and fpTEX, respectively.

You can always use

         M-x customize-group RET AUCTeX RET

in order to customize more stuff, or use the ‘Customize’ menu.

12. Load ‘preview/circ.tex’ into Emacs or XEmacs and see if you get the ‘Command’ menu. Try using it to LaTEX the file.

13. Check whether the ‘Preview’ menu is available in this file. Use it to generate previews for the document. If this barfs and tells you that image type ‘png’ is not supported, try adding the line

         (setq preview-image-type 'pnm)

at the end of your installed version of ‘preview-latex.el’. If this helps, complain to wherever you got your Emacs from: all current Emacs/XEmacs versions capable of running preview-latex by now can be compiled to support PNG images. Which is important, because PNM files take away vast amounts of disk space, and thus also of load/save time.

Well, that about is all. Have fun!

2.8 Customizing

Most of the site-specific customization should already have happened during configuration of AUCTEX. Any further customization can be done with customization buffers directly in Emacs. Just type M-x customize-group RET AUCTeX RET to open the customization group for AUCTEX or use the menu entries provided in the mode menus. Editing the file ‘tex-site.el’ as suggested in former versions of AUCTEX should not be done anymore because the installation routine will overwrite those changes.

You might check some variables with a special significance. They are accessible directly

by typing M-x customize-variable RET <variable> RET.
[User Option] TeX-macro-global

Directories containing the site’s TEX style files.

Normally, AUCTEX will only allow you to complete macros and environments which are built-in, specified in AUCTEX style files or defined by yourself. If you issue the M-x TeXauto-generate-global command after loading AUCTEX, you will be able to complete on all macros available in the standard style files used by your document. To do this, you must set this variable to a list of directories where the standard style files are located. The directories will be searched recursively, so there is no reason to list subdirectories explicitly. Automatic configuration will already have set the variable for you if it could use the program ‘kpsewhich’. In this case you normally don’t have to alter anything.

3 Quick Start

AUCTEX is a powerful program offering many features and configuration options. If you are new to AUCTEX this might be deterrent. Fortunately you do not have to learn everything at once. This Quick Start Guide will give you the knowledge of the most important commands and enable you to prepare your first LaTEX document with AUCTEX after only a few minutes of reading.

In this introduction, we assume that AUCTEX is already installed on your system. If this is not the case, you should read the file ‘INSTALL’ in the base directory of the unpacked distribution tarball. These installation instructions are available in this manual as well, Chapter 2 [Installation], page 5. We also assume that you are familiar with the way keystrokes are written in Emacs manuals. If not, have a look at the Emacs Tutorial in the Help menu.

If AUCTEX is installed, you might still need to activate it, by inserting

     (load "auctex.el" nil t t)

in your user init file.^{1} In order to get support for many of the LaTEX packages you will use in your documents, you should enable document parsing as well, which can be achieved by putting

     (setq TeX-auto-save t)
     (setq TeX-parse-self t)

into your init file. Finally, if you often use \include or \input, you should make AUCTEX aware of the multi-file document structure. You can do this by inserting

     (setq-default TeX-master nil)

into your init file. Each time you open a new file, AUCTEX will then ask you for a master file.

This Quick Start Guide covers two main topics: First we explain how AUCTEX helps you in editing your input file for TEX, LaTEX, and some other formats. Then we describe the functions that AUCTEX provides for processing the input files with LaTEX, BibTEX, etc., and for viewing and debugging.

3.1 Functions for editing TeX files

3.1.1 Making your TEX code more readable

AUCTEX can do syntax highlighting of your source code, that means commands will get special colors or fonts. You can enable it locally by typing M-x font-lock-mode RET. If you want to have font locking activated generally, enable global-font-lock-mode, e.g. with

M-x customize-variable RET global-font-lock-mode RET.

AUCTEX will indent new lines to indicate their syntactical relationship to the surrounding text. For example, the text of a \footnote or text inside of an environment will be indented relative to the text around it. If the indenting has gotten wrong after adding or deleting some characters, use ^{⟨}\underline{\overline{TAB}}^{⟩} to reindent the line, M-q for the whole paragraph, or

M-x LaTeX-fill-buffer RET for the whole buffer.

1 This usually is a file in your home directory called ‘.emacs’ if you are utilizing GNU Emacs or

‘.xemacs/init.el’ if you are using XEmacs.

3.1.2 Entering sectioning commands

Insertion of sectioning macros, that is ‘\chapter’, ‘\section’, ‘\subsection’, etc. and accompanying ‘\label’ commands may be eased by using C-c C-s. You will be asked for the section level. As nearly everywhere in AUCTEX, you can use the ^{⟨}\underline{\overline{TAB}}^{⟩} or ^{⟨}\underline{\overline{SPC}}^{⟩} key to get a list of available level names, and to auto-complete what you started typing. Next, you will be asked for the printed title of the section, and last you will be asked for a label to be associated with the section.

3.1.3 Inserting environments

Similarly, you can insert environments, that is ‘\begin{}’–‘\end{}’ pairs: Type C-c C-e, and select an environment type. Again, you can use ^{⟨}\underline{\overline{TAB}}^{⟩} or ^{⟨}\underline{\overline{SPC}}^{⟩} to get a list, and to complete what you type. Actually, the list will not only provide standard LaTEX environments, but also take your ‘\documentclass’ and ‘\usepackage’ commands into account if you have parsing enabled by setting TeX-parse-self to t. If you use a couple of environments frequently, you can use the up and down arrow keys (or M-p and M-n) in the minibuffer to get back to the previously inserted commands.

Some environments need additional arguments. Often, AUCTEX knows about this and asks you to enter a value.

3.1.4 Inserting macros

C-c C-m, or simply C-c RET will give you a prompt that asks you for a LaTEX macro. You can use ^{⟨}\underline{\overline{TAB}}^{⟩} for completion, or the up/down arrow keys (or M-p and M-n) to browse the command history. In many cases, AUCTEX knows which arguments a macro needs and will ask you for that. It even can differentiate between mandatory and optional arguments—for details, see Section 5.2 [Completion], page 29.

An additional help for inserting macros is provided by the possibility to complete macros right in the buffer. With point at the end of a partially written macro, you can complete it by typing M-TAB.

3.1.5 Changing the font

AUCTEX provides convenient keyboard shortcuts for inserting macros which specify the font to be used for typesetting certain parts of the text. They start with C-c C-f, and the last C- combination tells AUCTEX which font you want:

C-c C-f C-b

Insert bold face ‘\textbf{⋆}’ text.

C-c C-f C-i

Insert italics ‘\textit{⋆}’ text.

C-c C-f C-e

Insert emphasized ‘\emph{⋆}’ text.

C-c C-f C-s

Insert slanted ‘\textsl{⋆}’ text.

C-c C-f C-r

Insert roman \textrm{⋆} text.

C-c C-f C-f

Insert sans serif ‘\textsf{⋆}’ text.

C-c C-f C-t
          Insert typewriter ‘\texttt{⋆}’ text.
C-c C-f C-c

Insert small caps ‘\textsc{⋆}’ text.

C-c C-f C-d

Delete the innermost font specification containing point.

If you want to change font attributes of existing text, mark it as a region, and then invoke the commands. If no region is selected, the command will be inserted with empty braces, and you can start typing the changed text.

Most of those commands will also work in math mode, but then macros like \mathbf will be inserted.

3.1.6 Other useful features

AUCTEX also tries to help you when inserting the right “quote” signs for your language, dollar signs to typeset math, or pairs of braces. It offers shortcuts for commenting out text (C-c ; for the current region or C-c % for the paragraph you are in). The same keystrokes will remove the % signs, if the region or paragraph is commented out yet. With TeXfold-mode, you can hide certain parts (like footnotes, references etc.) that you do not edit currently. Support for Emacs’ outline mode is provided as well. And there’s more, but this is beyond the scope of this Quick Start Guide.

3.2 Creating and viewing output, debugging

3.2.1 One Command for LaTEX, helpers, viewers, and printing

If you have typed some text and want to run LaTEX (or TEX, or other programs—see below) on it, type C-c C-c. If applicable, you will be asked whether you want to save changes, and which program you want to invoke. In many cases, the choice that AUCTEX suggests will be just what you want: first latex, then a viewer. If a latex run produces or changes input files for makeindex, the next suggestion will be to run that program, and AUCTEX knows that you need to run latex again afterwards—the same holds for BibTEX.

When no processor invocation is necessary anymore, AUCTEX will suggest to run a viewer, or you can chose to create a PostScript file using dvips, or to directly print it.

At this place, a warning needs to be given: First, although AUCTEX is really good in detecting the standard situations when an additional latex run is necessary, it cannot detect it always. Second, the creation of PostScript files or direct printing currently only works when your output file is a DVI file, not a PDF file.

Ah, you didn’t know you can do both? That brings us to the next topic.

3.2.2 Choosing an output format

From a LaTEX file, you can produce DVI output, or a PDF file directly via pdflatex. You can switch on source specials for easier navigation in the output file, or tell latex to stop after an error (usually \noninteractive is used, to allow you to detect all errors in a single run).

These options are controlled by toggles, the keystrokes should be easy to memorize:

C-c C-t C-p

This command toggles between DVI and PDF output

C-c C-t C-i

toggles interactive mode

C-c C-t C-s

toggles source specials support

C-c C-t C-o

toggles usage of Omega/lambda.

3.2.3 Debugging LaTEX

When AUCTEX runs a program, it creates an output buffer in which it displays the output of the command. If there is a syntactical error in your file, latex will not complete successfully. AUCTEX will tell you that, and you can get to the place where the first error occured by pressing C-c ` (the last character is a backtick). The view will be split in two windows, the output will be displayed in the lower buffer, and both buffers will be centered around the place where the error ocurred. You can then try to fix it in the document buffer, and use the same keystrokes to get to the next error. This procedure may be repeated until all errors have been dealt with. By pressing C-c C-w (TeX-toggle-debug-boxes) you can toggle whether AUCTEX should notify you of overfull and underfull boxes in addition to regular errors.

If a command got stuck in a seemingly infinite loop, or you want to stop execution for other reasons, you can use C-c C-k (for “kill”). Similar to C-l, which centers the buffer you are in around your current position, C-c C-l centers the output buffer so that the last lines added at the bottom become visible.

3.2.4 Running LaTEX on parts of your document

If you want to check how some part of your text looks like, and do not want to wait until the whole document has been typeset, then mark it as a region and use C-c C-r. It behaves just like C-c C-c, but it only uses the document preamble and the region you marked.

If you are using \include or \input to structure your document, try C-c C-b while you are editing one of the included files. It will run latex only on the current buffer, using the preamble from the master file.

4 Inserting Frequently Used Commands

The most commonly used commands/macros of AUCTEX are those which simply insert templates for often used TEX and/or LaTEX/ConTEXt constructs, like font changes, handling of environments, etc. These features are very simple, and easy to learn, and help you avoid stupid mistakes like mismatched braces, or ‘\begin{}’-‘\end{}’ pairs.

4.1 Insertion of Quotes, Dollars, and Braces

In TEX, literal double quotes ‘"like this"’ are seldom used, instead two single quotes are used ‘``like this''’. To help you insert these efficiently, AUCTEX allows you to continue to press " to insert two single quotes. To get a literal double quote, press " twice.

[Command] TeX-insert-quote count

(") Insert the appropriate quote marks for TEX.

Inserts the value of TeX-open-quote (normally ‘``’) or TeX-close-quote (normally ‘''’) depending on the context. With prefix argument, always inserts ‘"’ characters.

[User Option] TeX-open-quote

String inserted by typing " to open a quotation.

[User Option] TeX-close-quote

String inserted by typing " to close a quotation.

[User Option] TeX-quote-after-quote

Determines the behavior of ". If it is non-nil, typing " will insert a literal double quote. The respective values of TeX-open-quote and TeX-close-quote will be inserted after typing " once again.

The ‘babel’ package provides special support for the requirements of typesetting quotation marks in many different languages. If you use this package, either directly or by loading a language-specific style file, you should also use the special commands for quote insertion instead of the standard quotes shown above. AUCTEX is able to recognize several of these languages and will change quote insertion accordingly. See Section 10.1 [European], page 55, for details about this feature and how to control it.

In case you are using the ‘csquotes’ package, you should customize LaTeX-csquotes-

open-quote, LaTeX-csquotes-close-quote and LaTeX-csquotes-quote-after-quote.

The quotation characters will only be used if both variables—LaTeX-csquotes-open-quote and LaTeX-csquotes-close-quote—are non-empty strings. But then the ‘csquotes’- related values will take precedence over the language-specific ones.

In AUCTEX, dollar signs should match like they do in TEX. This has been partially implemented, we assume dollar signs always match within a paragraph. The first ‘$’ you insert in a paragraph will do nothing special. The second ‘$’ will match the first. This will be indicated by moving the cursor temporarily over the first dollar sign. If you enter a dollar sign that matches a double dollar sign ‘$$’ AUCTEX will automatically insert two dollar signs. If you enter a second dollar sign that matches a single dollar sign, the single dollar sign will automatically be converted to a double dollar sign.

[Command] TeX-insert-dollar arg

($) Insert dollar sign.

Show matching dollar sign if this dollar sign end the TEX math mode. Ensure double dollar signs match up correctly by inserting extra dollar signs when needed. With optional arg, insert that many dollar signs.

To avoid unbalanced braces, it is useful to insert them pairwise. You can do this by typing C-c {.

[Command] TeX-insert-braces

(C-c {) Make a pair of braces and position the cursor to type inside of them. If there is an active region, put braces around it and leave point after the closing brace.

4.2 Inserting Font Specifiers

Perhaps the most used keyboard commands of AUCTEX are the short-cuts available for easy insertion of font changing macros.

If you give an argument (that is, type C-u) to the font command, the innermost font will be replaced, i.e. the font in the TEX group around point will be changed. The following table shows the available commands, with ⋆ indicating the position where the text will be inserted.

C-c C-f C-b

Insert bold face ‘\textbf{⋆}’ text.

C-c C-f C-i

Insert italics ‘\textit{⋆}’ text.

C-c C-f C-e

Insert emphasized ‘\emph{⋆}’ text.

C-c C-f C-s

Insert slanted ‘\textsl{⋆}’ text.

C-c C-f C-r

Insert roman \textrm{⋆} text.

C-c C-f C-f

Insert sans serif ‘\textsf{⋆}’ text.

C-c C-f C-t
          Insert typewriter ‘\texttt{⋆}’ text.
C-c C-f C-c

Insert small caps ‘\textsc{⋆}’ text.

C-c C-f C-d

Delete the innermost font specification containing point.

[Command] TeX-font arg

(C-c C-f) Insert template for font change command.

If replace is not nil, replace current font. what determines the font to use, as specified

     by TeX-font-list.

[User Option] TeX-font-list

List of fonts used by TeX-font.

Each entry is a list with three elements. The first element is the key to activate the font. The second element is the string to insert before point, and the third element is the string to insert after point. An optional fourth element means always replace if not nil.

4.3 Inserting chapters, sections, etc.

Insertion of sectioning macros, that is ‘\chapter’, ‘\section’, ‘\subsection’, etc. and accompanying ‘\label’’s may be eased by using C-c C-s. This command is highly customizable, the following describes the default behavior.

When invoking you will be asked for a section macro to insert. An appropriate default is automatically selected by AUCTEX, that is either: at the top of the document; the top level sectioning for that document style, and any other place: The same as the last occurring sectioning command.

Next, you will be asked for the actual name of that section, and last you will be asked for a label to be associated with that section. The label will be prefixed by the value specified

in LaTeX-section-hook.
[Command] LaTeX-section arg

(C-c C-s) Insert a sectioning command.

Determine the type of section to be inserted, by the argument arg.

• If arg is nil or missing, use the current level.

• If arg is a list (selected by C-u), go downward one level.

• If arg is negative, go up that many levels.

• If arg is positive or zero, use absolute level:

+ 0 : part + 1 : chapter + 2 : section + 3 : subsection + 4 : subsubsection + 5 : paragraph + 6 : subparagraph

The following variables can be set to customize the function.

     LaTeX-section-hook

Hooks to be run when inserting a section.

     LaTeX-section-label

Prefix to all section references.

The precise behavior of LaTeX-section is defined by the contents of LaTeX-section-

hook.

[User Option] LaTeX-section-hook

List of hooks to run when a new section is inserted. The following variables are set before the hooks are run

level Numeric section level, default set by prefix arg to LaTeX-section.

name Name of the sectioning command, derived from level.

title The title of the section, default to an empty string.

toc Entry for the table of contents list, default nil.

done-mark Position of point afterwards, default nil meaning after the inserted text.

A number of hooks are already defined. Most likely, you will be able to get the desired functionality by choosing from these hooks.

LaTeX-section-heading

Query the user about the name of the sectioning command. Modifies level and name.

LaTeX-section-title

Query the user about the title of the section. Modifies title.

LaTeX-section-toc

Query the user for the toc entry. Modifies toc.

LaTeX-section-section

Insert LaTEX section command according to name, title, and toc. If toc is nil, no toc entry is inserted. If toc or title are empty strings, done-mark will be placed at the point they should be inserted.

LaTeX-section-label

Insert a label after the section command. Controlled by the variable

          LaTeX-section-label.

To get a full featured LaTeX-section command, insert

     (setq LaTeX-section-hook
           '(LaTeX-section-heading
     LaTeX-section-title
     LaTeX-section-toc
     LaTeX-section-section
     LaTeX-section-label))

in your ‘.emacs’ file.

The behavior of LaTeX-section-label is determined by the variable LaTeX-section-

label.
[User Option] LaTeX-section-label

Default prefix when asking for a label. If it is a string, it is used unchanged for all kinds of sections. If it is nil, no label is inserted. If it is a list, the list is searched for a member whose car is equal to the name of the sectioning command being inserted. The cdr is then used as the prefix. If the name is not found, or if the cdr is nil, no label is inserted. By default, chapters have a prefix of ‘cha:’ while sections and subsections have a prefix of ‘sec:’. Labels are not automatically inserted for other types of sections.

4.4 Inserting Environment Templates

A large apparatus is available that supports insertions of environments, that is ‘\begin{}’ — ‘\end{}’ pairs.

AUCTEX is aware of most of the actual environments available in a specific document. This is achieved by examining your ‘\documentclass’ command, and consulting a precompiled list of environments available in a large number of styles.

You insert an environment with C-c C-e, and select an environment type. Depending on the environment, AUCTEX may ask more questions about the optional parts of the selected environment type. With C-u C-c C-e you will change the current environment.

[Command] LaTeX-environment arg

(C-c C-e) AUCTEX will prompt you for an environment to insert. At this prompt, you may press ^{⟨}\underline{\overline{TAB}}^{⟩} or ^{⟨}\underline{\overline{SPC}}^{⟩} to complete a partially written name, and/or to get a list of available environments. After selection of a specific environment AUCTEX may prompt you for further specifications. If the optional argument arg is not-nil (i.e. you have given a prefix argument), the current environment is modified and no new environment is inserted.

As a default selection, AUCTEX will suggest the environment last inserted or, as the first choice the value of the variable LaTeX-default-environment.

[User Option] LaTeX-default-environment

Default environment to insert when invoking ‘LaTeX-environment’ first time.

If the document is empty, or the cursor is placed at the top of the document, AUCTEX will default to insert a ‘document’ environment.

Most of these are described further in the following sections, and you may easily specify more. See Section 4.4.5 [Customizing environments], page 27.

You can close the current environment with C-c ], but we suggest that you use C-c C-e to insert complete environments instead.

[Command] LaTeX-close-environment

(C-c ]) Insert an ‘\end’ that matches the current environment.

4.4.1 Equations

When inserting equation-like environments, the ‘\label’ will have a default prefix, which is controlled by the following variables:

[User Option] LaTeX-equation-label

Prefix to use for ‘equation’ labels.

[User Option] LaTeX-eqnarray-label

Prefix to use for ‘eqnarray’ labels.

[User Option] LaTeX-amsmath-label

Prefix to use for amsmath equation labels. Amsmath equations include ‘align’,

     ‘alignat’, ‘xalignat’, ‘aligned’, ‘flalign’ and ‘gather’.

4.4.2 Floats

Figures and tables (i.e., floats) may also be inserted using AUCTEX. After choosing either ‘figure’ or ‘table’ in the environment list described above, you will be prompted for a number of additional things.

float position This is the optional argument of float environments that controls how they are placed in the final document. In LaTEX this is a sequence of the letters ‘htbp’ as described in the LaTEX manual. The value will default to the value of

          LaTeX-float.

caption This is the caption of the float. The default is to insert the caption at the bottom of the float. You can specify floats where the caption should be placed

          at the top with LaTeX-top-caption-list.

label The label of this float. The label will have a default prefix, which is controlled

          by the variables LaTeX-figure-label and LaTeX-table-label.

Moreover, you will be asked if you want the contents of the float environment to be horizontally centered. Upon a positive answer a ‘\centering’ macro will be inserted at the beginning of the float environment.

[User Option] LaTeX-float

Default placement for floats.

[User Option] LaTeX-figure-label

Prefix to use for figure labels.

[User Option] LaTeX-table-label

Prefix to use for table labels.

[User Option] LaTeX-top-caption-list

List of float environments with top caption.

4.4.3 Itemize-like

In an itemize-like environment, nodes (i.e., ‘\item’s) may be inserted using C-c ^{⟨}\underline{\overline{LFD}}^{⟩}.

[Command] LaTeX-insert-item

(C-c ^{⟨}\underline{\overline{LFD}}^{⟩}) Close the current item, move to the next line and insert an appropriate ‘\item’ for the current environment. That is, ‘itemize’ and ‘enumerate’ will have ‘\item ’ inserted, while ‘description’ will have ‘\item[]’ inserted.

4.4.4 Tabular-like

When inserting Tabular-like environments, that is, ‘tabular’ ‘array’ etc., you will be prompted for a template for that environment. Related variables:

[User Option] LaTeX-default-format

Default format string for array and tabular environments.

[User Option] LaTeX-default-position

Default position string for array and tabular environments. If nil, act like the empty string is given, but don’t prompt for a position.

4.4.5 Customizing environments

See Section 12.3 [Adding Environments], page 65, for how to customize the list of known environments.

5 Advanced Editing Features

The previous chapter described how to write the main body of the text easily and with a minimum of errors. In this chapter we will describe some features for entering more specialized sorts of text, for formatting the source by indenting and filling and for navigating through the document.

5.1 Entering Mathematics

TEX is written by a mathematician, and has always contained good support for formatting mathematical text. AUCTEX supports this tradition, by offering a special minor mode for entering text with many mathematical symbols. You can enter this mode by typing C-c ~.

[Command] LaTeX-math-mode

(C-c ~) Toggle LaTeX Math mode. This is a minor mode rebinding the key LaTeXmath-abbrev-prefix to allow easy typing of mathematical symbols. ` will read a character from the keyboard, and insert the symbol as specified in LaTeX-math-default and LaTeX-math-list. If given a prefix argument, the symbol will be surrounded by dollar signs.

You can use another prefix key (instead of `) by setting the variable LaTeX-math-

abbrev-prefix.

To enable LaTeX Math mode by default, add the following in your ‘.emacs’ file:

(add-hook 'LaTeX-mode-hook 'LaTeX-math-mode)
[User Option] LaTeX-math-abbrev-prefix

A string containing the prefix of LaTeX-math-mode commands; This value defaults to

`.

The string has to be a key or key sequence in a format understood by the kbd macro. This corresponds to the syntax usually used in the manuals for Emacs Emacs Lisp.

The variable LaTeX-math-list allows you to add your own mappings.

[User Option] LaTeX-math-list

A list containing user-defined keys and commands to be used in LaTeX Math mode. Each entry should be a list of two to four elements. First, the key to be used after LaTeX-math-abbrev-prefix for macro insertion. If it is nil, the symbol has no associated keystroke (it is available in the menu, though). Second, a string representing the name of the macro (without a leading backslash.) Third, a string representing the name of a submenu the command should be added to. Use a list of strings in case of nested menus. Fourth, the position of a Unicode character to be displayed in the menu alongside the macro name. This is an integer value.

[User Option] LaTeX-math-menu-unicode

Whether the LaTeX menu should try using Unicode for effect. Your Emacs built must be able to display include Unicode characters in menus for this feature.

AUCTEX’s reference card ‘tex-ref.tex’ includes a list of all math mode commands. AUCTEX can help you write subscripts and superscripts in math constructs by auto-

matically inserting a pair of braces after typing ^{⟨} ^{⟩} or ^{⟨}\underline{\overline{^}}^{⟩} respectively and putting point between the braces. In order to enable this feature, set the variable TeX-electric-sub-and-superscript to a non-nil value.

[User Option] TeX-electric-sub-and-superscript

If non-nil, insert braces after typing ^{⟨}\underline{\overline{^}}^{⟩} and ^{⟨} ^{⟩} in math mode.

5.2 Completion

Emacs lisp programmers probably know the lisp-complete-symbol command, usually bound to M-^{⟨}\underline{\overline{TAB}}^{⟩}. Users of the wonderful ispell mode know and love the ispell-complete-word command from that package. Similarly, AUCTEX has a TeX-complete-symbol command, usually bound to M-^{⟨}\underline{\overline{TAB}}^{⟩}. Using LaTeX-complete-symbol makes it easier to type and remember the names of long LaTEX macros.

In order to use TeX-complete-symbol, you should write a backslash and the start of the macro. Typing M-^{⟨}\underline{\overline{TAB}}^{⟩} will now complete as much of the macro, as it unambiguously can. For example, if you type “\renewc’’ and then M-^{⟨}\underline{\overline{TAB}}^{⟩}, it will expand to “\renewcommand’’.

[Command] TeX-complete-symbol

(M-^{⟨}\underline{\overline{TAB}}^{⟩}) Complete TEX symbol before point.

A more direct way to insert a macro is with TeX-insert-macro, bound to C-c C-m. It has the advantage over completion that it knows about the argument of most standard LaTEX macros, and will prompt for them. It also knows about the type of the arguments, so it will for example give completion for the argument to ‘\include’. Some examples are listed below.

[Command] TeX-insert-macro

(C-c C-m or C-c RET) Prompt (with completion) for the name of a TEX macro, and if AUCTEX knows the macro, prompt for each argument.

As a default selection, AUCTEX will suggest the macro last inserted or, as the first choice the value of the variable TeX-default-macro.

[User Option] TeX-insert-macro-default-style

Specifies whether TeX-insert-macro will ask for all optional arguments.

If set to the symbol show-optional-args, TeX-insert-macro asks for optional argu-

     ments of TEX macros. If set to mandatory-args-only, TeX-insert-macro asks only

for mandatory arguments. When TeX-insert-macro is called with prefix argument (C-u), it’s the other way round. Note that for some macros, there are special mechanisms, e.g. LaTeX-

     includegraphics-options-alist.
[User Option] TeX-default-macro

Default macro to insert when invoking TeX-insert-macro first time.

A faster alternative is to bind the function TeX-electric-macro to ‘\’. This can be done by setting the variable TeX-electric-escape

[User Option] TeX-electric-escape

If this is non-nil when AUCTEX is loaded, the TEX escape character ‘\’ will be bound

     to TeX-electric-macro

The difference between TeX-insert-macro and TeX-electric-macro is that space will complete and exit from the minibuffer in TeX-electric-macro. Use ^{⟨}\underline{\overline{TAB}}^{⟩} if you merely want to complete.

[Command] TeX-electric-macro

Prompt (with completion) for the name of a TEX macro, and if AUCTEX knows the macro, prompt for each argument. Space will complete and exit.

By default AUCTEX will put an empty set braces ‘{}’ after a macro with no arguments to stop it from eating the next whitespace. This can be stopped by entering LaTeX-math-mode, see Section 5.1 [Mathematics], page 28, or by setting TeX-insert-braces to nil.

[User Option] TeX-insert-braces

If non-nil, append a empty pair of braces after inserting a macro.

Completions work because AUCTEX can analyze TEX files, and store symbols in emacs lisp files for later retrieval. See Chapter 11 [Automatic], page 59, for more information.

AUCTEX will also make completion for many macro arguments, for example existing

labels when you enter a ‘\ref’ macro with TeX-insert-macro or TeX-electric-macro,

and BibTEX entries when you enter a ‘\cite’ macro. For this kind of completion to work, parsing must be enabled as described in see Chapter 9 [Parsing Files], page 53. For ‘\cite’ you must also make sure that the BibTEX files have been saved at least once after you enabled automatic parsing on save, and that the basename of the BibTEX file does not conflict with the basename of one of TEX files.

5.3 Commenting

It is often necessary to comment out temporarily a region of TEX or LaTEX code. This can be done with the commands C-c ; and C-c %. C-c ; will comment out all lines in the current region, while C-c % will comment out the current paragraph. Type C-c ; again to uncomment all lines of a commented region, or C-c % again to uncomment all comment lines around point. These commands will insert or remove a single ‘%’ respectively.

[Command] TeX-comment-or-uncomment-region

(C-c ;) Add or remove ‘%’ from the beginning of each line in the current region. Uncommenting works only if the region encloses solely commented lines. If AUCTEX should not try to guess if the region should be commented or uncommented the com-

     mands TeX-comment-region and TeX-uncomment-region can be used to explicitly

comment or uncomment the region in concern.

[Command] TeX-comment-or-uncomment-paragraph

(C-c %) Add or remove ‘%’ from the beginning of each line in the current paragraph. When removing ‘%’ characters the paragraph is considered to consist of all preceding and succeeding lines starting with a ‘%’, until the first non-comment line.

5.4 Indenting

Indentation means the addition of whitespace at the beginning of lines to reflect special syntactical constructs. This makes it easier to see the structure of the document, and to catch errors such as a missing closing brace. Thus, the indentation is done for precisely the same reasons that you would indent ordinary computer programs.

Indentation is done by LaTEX environments and by TEX groups, that is the body of an environment is indented by the value of LaTeX-indent-level (default 2). Also, items of an ‘itemize-like’ environment are indented by the value of LaTeX-item-indent, default −2. If more environments are nested, they are indented ‘accumulated’ just like most programming languages usually are seen indented in nested constructs.

You can explicitely indent single lines, usually by pressing ^{⟨}\underline{\overline{TAB}}^{⟩}, or marked regions by calling indent-region on it. If you have auto-fill-mode enabled and a line is broken while you type it, Emacs automatically cares about the indentation in the following line. If you want to have a similar behavior upon typing ^{⟨}\underline{\overline{RET}}^{⟩}, you can customize the variable TeX-newline-function and change the default of newline which does no indentation to

newline-and-indent which indents the new line or reindent-then-newline-and-indent

which indents both the current and the new line.

There are certain LaTEX environments which should be indented in a special way, like ‘tabular’ or ‘verbatim’. Those environments may be specified in the variable LaTeXindent-environment-list together with their special indentation functions. Taking the ‘verbatim’ environment as an example you can see that current-indentation is used as the indentation function. This will stop AUCTEX from doing any indentation in the environment if you hit ^{⟨}\underline{\overline{TAB}}^{⟩} for example.

There are environments in LaTeX-indent-environment-list which do not bring a special indentation function with them. This is due to the fact that first the respective functions are not implemented yet and second that filling will be disabled for the specified environments. This shall prevent the source code from being messed up by accidently filling those environments with the standard filling routine. If you think that providing special filling routines for such environments would be an appropriate and challenging task for you, you are invited to contribute. (See Section 5.5 [Filling], page 32, for further information about the filling functionality)

The check for the indentation function may be enabled or disabled by customizing the

variable LaTeX-indent-environment-check.

As a side note with regard to formatting special environments: Newer Emacsen include ‘align.el’ and therefore provide some support for formatting ‘tabular’ and ‘tabbing’ environments with the function align-current which will nicely align columns in the source code.

AUCTEX is able to format commented parts of your code just as any other part. This means LaTEX environments and TEX groups in comments will be indented syntactically correct if the variable LaTeX-syntactic-comments is set to t. If you disable it, comments will be filled like normal text and no syntactic indentation will be done.

Following you will find a list of most commands and variables related to indenting with a small summary in each case:

⟨\underline{\overline{TAB}}^{⟩} LaTeX-indent-line will indent the current line.

⟨\underline{\overline{LFD}}^{⟩} newline-and-indent inserts a new line (much like ^{⟨}\underline{\overline{RET}}^{⟩}) and moves the cursor to an appropriate position by the left margin. Most keyboards nowadays don’t have a linefeed key and C-j is tedious to type. Therefore you can customize AUCTEX to perform indentation (or to make coffee) upon typing ^{⟨}\underline{\overline{RET}}^{⟩} as well. The respective option is called TeX-newline-

     function.

C-j Alias for ^{⟨}\underline{\overline{LFD}}^{⟩}

[User Option] LaTeX-indent-environment-list

List of environments with special indentation. The second element in each entry is the function to calculate the indentation level in columns. The filling code currently cannot handle tabular-like environments which will be completely messed-up if you try to format them. This is why most of these environments are included in this customization option without a special indentation function. This will prevent that they get filled.

[User Option] LaTeX-indent-level

Number of spaces to add to the indentation for each ‘\begin’ not matched by a ‘\end’.

[User Option] LaTeX-item-indent

Number of spaces to add to the indentation for ‘\item’’s in list environments.

[User Option] TeX-brace-indent-level

Number of spaces to add to the indentation for each ‘{’ not matched by a ‘}’.

[User Option] LaTeX-syntactic-comments

If non-nil comments will be filled and indented according to LaTEX syntax. Otherwise they will be filled like normal text.

[User Option] TeX-newline-function

Used to specify the function which is called when ^{⟨}\underline{\overline{RET}}^{⟩} is pressed. This will normally be newline which simply inserts a new line. In case you want to have AUCTEX do indentation as well when you press ^{⟨}\underline{\overline{RET}}^{⟩}, use the built-in functions newline-and-

indent or reindent-then-newline-and-indent. The former inserts a new line and

indents the following line, i.e. it moves the cursor to the right position and therefore acts as if you pressed ^{⟨}\underline{\overline{LFD}}^{⟩}. The latter function additionally indents the current line. If you choose ‘Other’, you can specify your own fancy function to be called when ⟨\underline{\overline{RET}}^{⟩} is pressed.

5.5 Filling

Filling deals with the insertion of line breaks to prevent lines from becoming wider than what is specified in fill-column. The linebreaks will be inserted automatically if auto-fill-mode is enabled. In this case the source is not only filled but also indented automatically as you write it.

auto-fill-mode can be enabled for AUCTEX by calling turn-on-auto-fill in one of the hooks AUCTEX is running. For all text modes with text-mode-hook, for all AUCTEX

modes with TeX-mode-hook or for specific modes with plain-TeX-mode-hook, LaTeX-mode-
hook, ConTeXt-mode-hook or docTeX-mode-hook. As an example, if you want to enable

auto-fill-mode in LaTeX-mode, put the following into your init file:

     (add-hook 'LaTeX-mode-hook 'turn-on-auto-fill)

You can manually fill explicitely marked regions, paragraphs, environments, complete sections, or the whole buffer. (Note that manual filling in AUCTEX will indent the start of the region to be filled in contrast to many other Emacs modes.)

There are some syntactical constructs which are handled specially with regard to filling. These are so-called code comments and paragraph commands.

Code comments are comments preceded by code or text in the same line. Upon filling a region, code comments themselves will not get filled. Filling is done from the start of the region to the line with the code comment and continues after it. In order to prevent overfull lines in the source code, a linebreak will be inserted before the last non-comment word by default. This can be changed by customizing LaTeX-fill-break-before-code-comments. If you have overfull lines with code comments you can fill those explicitely by calling LaTeX-fill-paragraph or pressing M-q with the cursor positioned on them. This will add linebreaks in the comment and indent subsequent comment lines to the column of the comment in the first line of the code comment. In this special case M-q only acts on the current line and not on the whole paragraph.

Lines with ‘\par’ are treated similarly to code comments, i.e. ‘\par’ will be treated as paragraph boundary which should not be followed by other code or text. But it is not treated as a real paragraph boundary like an empty line where filling a paragraph would stop.

Paragraph commands like ‘\section’ or ‘\noindent’ (the list of commands is defined by LaTeX-paragraph-commands) are often to be placed in their own line(s). This means they should not be consecuted with any preceding or following adjacent lines of text. AUCTEX will prevent this from happening if you do not put any text except another macro after the end of the last brace of the respective macro. If there is other text after the macro, AUCTEX regards this as a sign that the macro is part of the following paragraph.

Here are some examples:

     \begin{quote}
       text text text text
     \begin{quote}\label{foo}
       text text text text

If you press M-q on the first line in both examples, nothing will change. But if you write

     \begin{quote} text
       text text text text

and press M-q, you will get

     \begin{quote} text text text text text

Besides code comments and paragraph commands, another speciality of filling in AUCTEX involves commented lines. You should be aware that these comments are treated as islands in the rest of the LaTEX code if syntactic filling is enabled. This means, for example, if you try to fill an environment with LaTeX-fill-environment and have the cursor placed on a commented line which does not have a surrounding environment inside the comment, AUCTEX will report an error.

The relevant commands and variables with regard to filling are:

C-c C-q C-p

LaTeX-fill-paragraph will fill and indent the current paragraph.

M-q  Alias for C-c C-q C-p
C-c C-q C-e

LaTeX-fill-environment will fill and indent the current environment. This may e.g. be the ‘document’ environment, in which case the entire document will be formatted.

C-c C-q C-s

LaTeX-fill-section will fill and indent the current logical sectional unit.

C-c C-q C-r

LaTeX-fill-region will fill and indent the current region.

[User Option] LaTeX-fill-break-at-separators

List of separators before or after which respectively linebreaks will be inserted if they do not fit into one line. The separators can be curly braces, brackets, switches for inline math (‘$’, ‘\(’, ‘\)’) and switches for display math (‘\[’, ‘\]’). Such formatting can be useful to make macros and math more visible or to prevent overfull lines in the LaTEX source in case a package for displaying formatted TEX output inside the Emacs buffer, like preview-latex, is used.

[User Option] LaTeX-fill-break-before-code-comments

Code comments are comments preceded by some other text in the same line. When a paragraph containing such a comment is to be filled, the comment start will be seen as a border after which no line breaks will be inserted in the same line. If the

option LaTeX-fill-break-before-code-comments is enabled (which is the default)

and the comment does not fit into the line, a line break will be inserted before the last non-comment word to minimize the chance that the line becomes overfull.

6 Controlling Screen Display

It is often desirable to get visual help of what markup code in a text actually does whithout having to decipher it explicitely. For this purpose Emacs and AUCTEX provide font locking (also known as syntax highlighting) which visually sets off markup code like macros or environments by using different colors or fonts. For example text to be typeset in italics can be displayed with an italic font in the editor as well, or labels and references get their own distinct color.

While font locking helps you grasp the purpose of markup code and separate markup from content, the markup code can still be distracting. AUCTEX lets you hide those parts and show them again at request with its built-in support for hiding macros and environments which we call folding here.

Besides folding of macros and environments, AUCTEX provides support for Emacs’ outline mode which lets you narrow the buffer content to certain sections of your text by hiding the parts not belonging to these sections.

6.1 Font Locking

Font locking is supposed to improve readability of the source code by highlighting certain keywords with different colors or fonts. It thereby lets you recognize the function of markup code to a certain extent without having to read the markup command. For general information on controlling font locking with Emacs’ Font Lock mode, see section “Font Lock Mode” in GNU Emacs Manual.

[User Option] TeX-install-font-lock

Once font locking is enabled globally or for the major modes provided by AUCTEX, the font locking patterns and functionality of font-latex are activated by default. You can switch to a different font locking scheme or disable font locking in AUCTEX by customizing the variable TeX-install-font-lock. Besides font-latex AUCTEX ships with a scheme which is derived from Emacs’ default LaTEX mode and activated by choosing tex-font-setup. Be aware that this scheme is not coupled with AUCTEX’s style system and not the focus of development. Therefore and due to font-latex being much more feature-rich the following explanations will only cover font-latex. In case you want to hook in your own fontification scheme, you can choose other and insert the name of the function which sets up your font locking patterns. If you want to disable fontification in AUCTEX completely, choose ignore.

font-latex provides many options for customization which are accessible with M-x customize-group RET font-latex RET. For this description the various options are explained in conceptional groups.

Macros

Highlighting of macros can be customized by adapting keyword lists which can be found in the customization group font-latex-keywords. The lists contain names of macros without the leading backslash.

Three types of macros can be handled differently with respect to fontification:

1. Commands of the form ‘\foo[bar]{baz}’ which consist of the macro itself, optional arguments in square brackets and mandatory arguments in curly braces. For the command itself the face font-lock-keyword-face will be used and for the optional arguments the face font-lock-variable-name-face. The face applied to the mandatory argument depends on the macro class represented by the respective built-in variables.

2. Declaration macros of the form ‘{\foo text}’ which consist of the macro which may be enclosed in a TEX group together with text to be affected by the macro. In case a TEX group is present, the macro will get the face font-lock-keyword-face and the text will get the face configured for the respective macro class. If no TEX group is present, the latter face will be applied to the macro itself.

3. Simple macros of the form ‘\foo’ which do not have any arguments or groupings. The respective face will be applied to the macro itself.

General macro classes

font-latex provides keyword lists for different macro classes which are described in the following table:

font-latex-match-function-keywords

Keywords for macros defining or related to functions, like ‘\newcommand’.

Type: ‘\macro[...]{...}’
Face: font-lock-function-name-face
font-latex-match-reference-keywords

Keywords for macros defining or related to references, like ‘\ref’.

Type: ‘\macro[...]{...}’
Face: font-lock-constant-face
font-latex-match-textual-keywords

Keywords for macros specifying textual content, like ‘\caption’.

Type: ‘\macro[...]{...}’
Face: font-lock-type-face
font-latex-match-variable-keywords

Keywords for macros defining or related to variables, like ‘\setlength’.

Type: ‘\macro[...]{...}{...}’
Face: font-lock-variable-name-face
font-latex-match-warning-keywords

Keywords for important macros, e.g. affecting line or page break, like

‘\clearpage’.

Type: ‘\macro’

Face: font-latex-warning-face

Sectioning commands

Sectioning commands are macros like ‘\chapter’ or ‘\section’. For these commands there are two fontification schemes which may be selected by customizing the variable font-

latex-fontify-sectioning.

[User Option] font-latex-fontify-sectioning

Per default sectioning commands will be shown in a larger, proportional font, which corresponds to a number for this variable. The font size varies with the sectioning level, e.g. ‘\part’ (font-latex-sectioning-0-face) has a larger font than

     ‘\paragraph’ (font-latex-sectioning-5-face). Typically, values from 1.05 to 1.3

for font-latex-fontify-sectioning give best results, depending on your font setup. If you rather like to use the base font and a different color, set the variable to the symbol ‘color’. In this case the face font-lock-type-face will be used to fontify the argument of the sectioning commands.

You can make font-latex aware of your own sectioning commands be adding them to the

keyword lists: font-latex-match-sectioning-0-keywords (font-latex-sectioning-
0-face) . . . font-latex-match-sectioning-5-keywords (font-latex-sectioning-5-
face).

Related to sectioning there is special support for slide titles which may be fontified with the face font-latex-slide-title-face. You can add macros which should appear in this

face by customizing the variable font-latex-match-slide-title-keywords.

Commands for changing fonts

LaTEX provides various macros for changing fonts or font attributes. For example, you can select an italic font with ‘\textit{...}’ or bold with ‘\textbf{...}’. An alternative way to specify these fonts is to use special macros in TEX groups, like ‘{\itshape ...}’ for italics and ‘{\bfseries ...}’ for bold. As mentioned above, we call the former variants commands and the latter declarations.

Besides the macros for changing fonts provided by LaTEX there is an infinite number of other macros—either defined by yourself for logical markup or defined by macro packages— which affect the font in the typeset text. While LaTEX’s built-in macros and macros of packages known by AUCTEX are already handled by font-latex, different keyword lists per type style and macro type are provided for entering your own macros which are listed in the table below.

font-latex-match-bold-command-keywords

Keywords for commands specifying a bold type style.

          Face: font-latex-bold-face
font-latex-match-italic-command-keywords

Keywords for commands specifying an italic font.

          Face: font-latex-italic-face
font-latex-match-math-command-keywords

Keywords for commands specifying a math font.

          Face: font-latex-math-face
font-latex-match-type-command-keywords

Keywords for commands specifying a typewriter font.

          Face: font-lock-type-face
font-latex-match-bold-declaration-keywords

Keywords for declarations specifying a bold type style.

          Face: font-latex-bold-face

font-latex-match-italic-declaration-keywords

Keywords for declarations specifying an italic font.

          Face: font-latex-italic-face
font-latex-match-type-declaration-keywords

Keywords for declarations specifying a typewriter font.

          Face: font-latex-type-face

Deactivating defaults of built-in keyword classes

font-latex ships with predefined lists of keywords for the classes described above. You can disable these defaults per class by customizing the variable font-latex-deactivated-keyword-classes. This is a list of strings for keyword classes to be deactivated. Valid entries are \"warning\", \"variable\", \"reference\", \"function\" , \"sectioning-0\", \"sectioning-1\", \"sectioning-2\", \"sectioning-3\", \"sectioning-4\", \"sectioning-5\", \"textual\", \"bold-command\", \"italic-command\", \"math-command\", \"typecommand\", \"bold-declaration\", \"italic-declaration\", \"type-declaration\".

You can also get rid of certain keywords only. For example if you want to remove highlighting of footnotes as references you can put the following stanza into your init file:

     (eval-after-load "font-latex"
       '(setq-default
         font-latex-match-reference-keywords-local
         (remove "footnote" font-latex-match-reference-keywords-local)))

But note that this means fiddling with font-latex’s internals and is not guaranteed to work in future versions of font-latex.

User-defined keyword classes

In case the customization options explained above do not suffice for your needs, you can specify your own keyword classes by customizing the variable font-latex-user-keyword-

classes.
[User Option] font-latex-user-keyword-classes

Every keyword class consists of four parts, a name, a list of keywords, a face and a specifier for the type of macros to be highlighted.

When adding new entries, you have to use unique values for the class names, i.e. they must not clash with names of the built-in keyword classes or other names given by you. Additionally the names must not contain spaces.

The keywords are names of commands you want to match omitting the leading backslash.

The face argument can either be an existing face or font specifications made by you. (The latter option is not available on XEmacs.)

There are three alternatives for the type of keywords—“Command with arguments”, “Declaration inside TEX group” and “Command without arguments”—which correspond with the macro types explained above.

Quotes

Text in quotation marks is displayed with the face font-latex-string-face. Besides the various forms of opening and closing double and single quotation marks, so-called guillemets (<<, >>) can be used for quoting. Because there are two styles of using them—French style: << text >>; German style: >>text<<—you can customize the variable font-latex-quotes to tell font-latex which type you are using if the correct value cannot be derived from document properties.

[User Option] font-latex-quotes

The default value of font-latex-quotes is ‘auto’ which means that font-latex will try to derive the correct type of quotation mark matching from document properties like the language option supplied to the babel LaTEX package. If the automatic detection fails for you and you mostly use one specific style you can set it to a specific language-dependent value as well. Set the value to ‘german’ if you are using >>German quotes<< and to ‘french’ if you are using << French quotes >>. font-latex will recognize the different ways these quotes can be given in your source code, i.e. (‘"<’, ‘">’), (‘<<’, ‘>>’) and the respective 8-bit variants.

Subscript and superscript in math

In order to make math constructs more readable, font-latex displays subscript and superscript parts in a smaller font and raised or lowered respectively. This fontification feature can

be controlled with the variables font-latex-fontify-script and font-latex-script-
display.
[User Option] font-latex-fontify-script

If non-nil, fontify subscript and superscript strings.

Note that this feature is not available on XEmacs, for which it is disabled per default. In GNU Emacs raising and lowering is not enabled for versions 21.3 and before due to it working not properly.

[User Option] font-latex-script-display

Display specification for subscript and superscript content. The car is used for subscript, the cdr is used for superscript. The feature is implemented using so-called display properties. For information on what exactly to specify for the values, see section “Other Display Specifications” in GNU Emacs Lisp Reference Manual.

Verbatim macros and environments

Usually it is not desirable to have content to be typeset verbatim highlighted according to LaTEX syntax. Therefore this content will be fontified uniformly with the face font-latex-

verbatim-face.

font-latex differentiates three different types of verbatim constructs for fontification. Macros with special characters like | as delimiters, macros with braces, and environments. Which macros and environments are recognized is controlled by the

variables LaTeX-verbatim-macros-with-delims, LaTeX-verbatim-macros-with-braces,
and LaTeX-verbatim-environments respectively.

Multi-line fontification

Font locking in LaTEX source code often involves constructs spanning more than one line of text. For these constructs to be handled correctly GNU Emacs as well as font-latex provide mechanisms for multi-line fontification which can be controlled by the variable font-latex-

do-multi-line.
[User Option] font-latex-do-multi-line

Control multi-line fontification.

Setting the variable to t will enable font-latex’s mechanism, setting it to nil will disable it. Setting it to ‘try-font-lock’ (the default) will use font-lock’s mechanism if available and font-latex’s method if not. Setting this variable will only have effect after resetting buffers controlled by font-latex or restarting Emacs.

Faces

In case you want to change the colors and fonts used by font-latex please refer to the faces mentioned in the explanations above and use M-x customize-face RET <face> RET. All faces defined by font-latex are accessible through a customization group by typing M-x

customize-group RET font-latex-highlighting-faces RET.

6.2 Folding Macros and Environments

A popular complaint about markup languages like TEX and LaTEX is that there is too much clutter in the source text and that one cannot focus well on the content. There are macros where you are only interested in the content they are enclosing, like font specifiers where the content might already be fontified in a special way by font locking. Or macros the content of which you only want to see when actually editing it, like footnotes or citations. Similarly you might find certain environments or comments distracting when trying to concentrate on the body of your document.

With AUCTEX’s folding functionality you can collapse those items and replace them by either a fixed string or the content of one of their arguments instead. If you want to make the original text visible again in order to view or edit it, move point sideways onto the placeholder (also called display string) or left-click with the mouse pointer on it. (The latter is currently only supported on Emacs.) The macro or environment will unfold automatically, stay open as long as point is inside of it and collapse again once you move point out of it. (Note that folding of environments currently does not work in every AUCTEX mode.)

In order to use this feature, you have to activate TeX-fold-mode which will activate the auto-reveal feature and the necessary commands to hide and show macros and environments. You can activate the mode in a certain buffer by typing the command M-x TeX-fold-mode RET or using the keyboard shortcut C-c C-o C-f. If you want to use it every time you edit a LaTEX document, add it to a hook:

     (add-hook 'LaTeX-mode-hook (lambda ()
                                  (TeX-fold-mode 1)))

If it should be activated in all AUCTEX modes, use TeX-mode-hook instead of LaTeX-

mode-hook.

Once the mode is active there are several commands available to hide and show macros, environments and comments:

[Command] TeX-fold-buffer

(C-c C-o C-b) Hide all foldable items in the current buffer according to the setting of TeX-fold-type-list. This command can also be used to refresh the whole buffer and hide any new macros and environments which were inserted after the last invocation of the command.

[User Option] TeX-fold-type-list

List of symbols determining the item classes to consider for folding. This can be macros, environments and comments. Per default only macros and environments are folded.

[User Option] TeX-fold-force-fontify

In order for all folded content to get the right faces, the whole buffer has to be fontified before folding is carried out. TeX-fold-buffer therefore will force fontification of unfontified regions. As this will prolong the time folding takes, you can prevent forced fontification by customizing the variable TeX-fold-force-fontify.

[User Option] TeX-fold-preserve-comments

By default items found in comments will be folded. If your comments often contain unfinished code this might lead to problems. Give this variable a non-nil value and foldable items in your comments will be left alone.

[Command] TeX-fold-region

(C-c C-o C-r) Hide all configured macros in the marked region.

[Command] TeX-fold-paragraph

(C-c C-o C-p) Hide all configured macros in the paragraph containing point.

[Command] TeX-fold-macro

(C-c C-o C-m) Hide the macro on which point currently is located. If the name of the macro is found in TeX-fold-macro-spec-list, the respective display string will be shown instead. If it is not found, the name of the macro in sqare brackets or the de-

fault string for unspecified macros (TeX-fold-unspec-macro-display-string) will

be shown, depending on the value of the variable TeX-fold-unspec-use-name.

[Command] TeX-fold-env

(C-c C-o C-e) Hide the environment on which point currently is located. The behavior regarding the display string is analogous to TeX-fold-macro and determined

by the variables TeX-fold-env-spec-list and TeX-fold-unspec-env-display-

string respectively.

[Command] TeX-fold-comment

(C-c C-o C-c) Hide the comment point is located on.

[Command] TeX-fold-clearout-buffer

(C-c C-o b) Permanently unfold all macros and environments in the current buffer.

[Command] TeX-fold-clearout-region

(C-c C-o r) Permanently unfold all macros and environments in the marked region.

[Command] TeX-fold-clearout-paragraph

(C-c C-o p) Permanently unfold all macros and environments in the paragraph containing point.

[Command] TeX-fold-clearout-item

(C-c C-o i) Permanently show the macro or environment on which point currently is located. In contrast to temporarily opening the macro when point is moved sideways onto it, the macro will be permanently unfolded and will not collapse again once point is leaving it.

[Command] TeX-fold-dwim

(C-c C-o C-o) Hide or show items according to the current context. If there is folded content, unfold it. If there is a marked region, fold all configured content in this region. If there is no folded content but a macro or environment, fold it.

The commands above will only take macros or environments into consideration which

are specified in the variable TeX-fold-macro-spec-list or TeX-fold-env-spec-list re-

spectively.

[User Option] TeX-fold-macro-spec-list

List of display strings or argument numbers and macros to fold. If you specify a number, the content of the first mandatory argument of a LaTEX macro will be used as the placeholder. The placeholder is made by copying the text from the buffer together with its properties, i.e. its face as well. If fontification has not happened when this is done (e.g. because of lazy font locking) the intended fontification will not show up. As a workaround you can leave Emacs idle a few seconds and wait for stealth font locking to finish before you fold the buffer. Or you just re-fold the buffer with TeX-fold-buffer when you notice a wrong fontification.

[User Option] TeX-fold-env-spec-list

List of display strings or argument numbers and environments to fold. Argument numbers refer to the ‘\begin’ statement. That means if you have e.g.

‘\begin{tabularx}{\linewidth}{XXX} ... \end{tabularx}’ and specify 3 as the

argument number, the resulting display string will be “XXX”.

[User Option] TeX-fold-unspec-macro-display-string

Default display string for macros which are not specified in TeX-fold-macro-spec-

list.
[User Option] TeX-fold-unspec-env-display-string

Default display string for environments which are not specified in TeX-fold-env-

spec-list.
[User Option] TeX-fold-unspec-use-name

If non-nil the name of the macro or environment surrounded by square brackets is used as display string, otherwise the defaults specified in TeX-fold-unspec-macro-

display-string or TeX-fold-unspec-env-display-string respectively.

When you hover with the mouse pointer over folded content, its original text will be shown in a tooltip or the echo area depending on Tooltip mode being activate. In order to avoid exorbitantly big tooltips and to cater for the limited space in the echo area the content will be cropped after a certain amount of characters defined by the variable TeX-

fold-help-echo-max-length.
[User Option] TeX-fold-help-echo-max-length

Maximum length of original text displayed in a tooltip or the echo area for folded content. Set it to zero in order to disable this feature.

6.3 Outlining the Document

AUCTEX supports the standard outline minor mode using LaTEX/ConTEXt sectioning commands as header lines. See section “Outline Mode” in GNU Emacs Manual.

You can add your own headings by setting the variable TeX-outline-extra.

[Variable] TeX-outline-extra

List of extra TEX outline levels. Each element is a list with two entries. The first entry is the regular expression matching a header, and the second is the level of the header. A ‘^’ is automatically prepended to the regular expressions in the list, so they must match text at the beginning of the line.

See LaTeX-section-list or ConTeXt-INTERFACE-section-list for existing header

levels.

The following example add ‘\item’ and ‘\bibliography’ headers, with ‘\bibliography’ at the same outline level as ‘\section’, and ‘\item’ being below ‘\subparagraph’.

(setq TeX-outline-extra
      '(("[ \t]*\\\\\\(bib\\)?item\\b" 7)
("\\\\bibliography\\b" 2)))

You may want to check out the unbundled ‘out-xtra’ package for even better outline support. It is available from your favorite emacs lisp archive.

7 Starting Processors, Viewers and Other Programs

The most powerful features of AUCTEX may be those allowing you to run (La)TEX/ConTEXt and other external commands like BibTEX and makeindex from within Emacs, viewing and printing the results, and moreover allowing you to debug your documents.

AUCTEX comes with a special tool bar for TEX and LaTEX which provides buttons for the most important commands. You can enable or disable it by customizing the options

plain-TeX-enable-toolbar and LaTeX-enable-toolbar in the TeX-tool-bar customiza-

tion group.

7.1 Executing Commands

Formatting the document with TEX, LaTEX or ConTEXt, viewing with a previewer, printing the document, running BibTEX, making an index, or checking the document with lacheck or chktex all require running an external command.

There are two ways to run an external command, you can either run it on all of the current documents with TeX-command-master, or on the current region with TeX-command-region. A special case of running TEX on a region is TeX-command-buffer which differs from TeX-command-master if the current buffer is not its own master file.

[Command] TeX-command-master

(C-c C-c) Query the user for a command, and run it on the master file associated with the current buffer. The name of the master file is controlled by the variable TeXmaster. The available commands are controlled by the variable TeX-command-list.

See Chapter 2 [Installation], page 5, for a discussion about TeX-command-list and Chapter 8 [Multifile], page 51 for a discussion about TeX-master.

[Command] TeX-command-region

(C-c C-r) Query the user for a command, and run it on the “region file”. Some commands (typically those invoking TEX or LaTEX) will write the current region into the region file, after extracting the header and tailer from the master file. If mark is inactive (which can happen with transient-mark-mode), use the old region. The name of the region file is controlled by the variable TeX-region. The name of the master file is controlled by the variable TeX-master. The header is all text up to the line matching the regular expression TeX-header-end. The trailer is all text from the line matching the regular expression TeX-trailer-start. The available commands are controlled by the variable TeX-command-list.

[Command] TeX-pin-region

(C-c C-t C-r) If you don’t have a mode like transient-mark-mode active, where marks get disabled automatically, the region would need to get properly set before each call to TeX-command-region. If you fix the current region with C-c C-t C-r, then it will get used for more commands even though mark and point may change. An explicitly activated mark, however, will always define a new region when calling

TeX-command-region.

[Command] TeX-command-buffer

(C-c C-b) Query the user for a command, and run it on the “region file”. Some commands (typically those invoking TEX or LaTEX) will write the current buffer into the region file, after extracting the header and tailer from the master file. See above for details.

AUCTEX will allow one process for each document, plus one process for the region file to be active at the same time. Thus, if you are editing n different documents, you can have n plus one processes running at the same time. If the last process you started was on the region, the commands described in Section 7.3 [Debugging], page 48 and Section 7.5 [Control], page 49 will work on that process, otherwise they will work on the process associated with the current document.

[User Option] TeX-region

The name of the file for temporarily storing the text when formatting the current region.

[User Option] TeX-header-end

A regular expression matching the end of the header. By default, this is

     ‘\begin{document}’ in LaTEX mode and ‘%**end of header’ in TEX mode.
[User Option] TeX-trailer-start

A regular expression matching the start of the trailer. By default, this is ‘\end{document}’ in LaTEX mode and ‘\bye’ in TEX mode.

AUCTEX will try to guess what command you want to invoke, but by default it will assume that you want to run TEX in TEX mode and LaTEX in LaTEX mode. You can overwrite this by setting the variable TeX-command-default.

[User Option] TeX-command-default

The default command to run in this buffer. Must be an entry in TeX-command-list.

   If you want to overwrite the values of TeX-header-end, TeX-trailer-start, or TeX-

command-default, you can do that for all files by setting them in either TeX-mode-hook, plain-TeX-mode-hook, or LaTeX-mode-hook. To overwrite them for a single file, define them as file variables (see section “File Variables” in The Emacs Editor). You do this by putting special formatted text near the end of the file.

     %%% Local Variables:
     %%% TeX-header-end: "% End-Of-Header"
     %%% TeX-trailer-start: "% Start-Of-Trailer"
     %%% TeX-command-default: "SliTeX"
     %%% End:

AUCTEX will try to save any buffers related to the document, and check if the document needs to be reformatted. If the variable TeX-save-query is non-nil, AUCTEX will query before saving each file. By default AUCTEX will check emacs buffers associated with files in the current directory, in one of the TeX-macro-private directories, and in the TeX-macro-global directories. You can change this by setting the variable TeX-check-path.

[User Option] TeX-check-path

Directory path to search for dependencies.

If nil, just check the current file. Used when checking if any files have changed.

[Command] TeX-PDF-mode

(C-c C-t C-p) This command toggles the PDF mode of AUCTEX, a buffer-local minor mode. You can customize TeX-PDF-mode to give it a different default. The default is used when AUCTEX does not have additional clue about what a document might want. This option usually results in calling either PDFTEX or ordinary TEX.

[User Option] TeX-DVI-via-PDFTeX

If this is set, DVI will also be produced by calling PDFTEX, setting \pdfoutput=0. This makes it possible to use packages like ‘pdfcprot’ even when producing DVI files. Some modern TEX distributions, e.g. teTEX 3.0, do this anyway, so that you need not enable it within AUCTEX.

[Command] TeX-interactive-mode

(C-c C-t C-i) This command toggles the interactive mode of AUCTEX, a global minor mode. You can customize TeX-interactive-mode to give it a different default. In interactive mode, TEX will pause with an error prompt when errors are encountered and wait for the user to type something.

[Command] TeX-source-specials-mode

(C-c C-t C-s) toggles Source Special support. Source Specials will move the DVI viewer to the location corresponding to point (forward search), and it will use ‘emacsclient’ or ‘gnuclient’ to have the previewer move Emacs to a location corresponding to a control-click in the previewer window. See Section 7.2 [Viewing], page 46.

You can permanently activate TeX-source-specials-mode with

     (TeX-source-specials-mode 1)

or by customizing the variable TeX-source-specials-mode. There is a bunch of customization options, use customize-group on the group ‘TeX-source-specials’ to find out more.

It has to be stressed very strongly however, that Source Specials can cause differences in page breaks, in spacing, can seriously interfere with various packages and should thus never be used for the final version of a document. In particular, fine-tuning the page breaks should be done with Source Specials switched off.

[Command] TeX-Omega-mode

(C-c C-t C-o) This command toggles the use of the Omega (Ω) mode of AUCTEX, a buffer-local minor mode. If it is switched on, omega will be used instead of tex, and

lambda instead of latex.

7.2 Viewing the formatted output

7.2.1 Starting viewers and customizing their invocation

AUCTEX allows you to start external programs for previewing your document. These are normally invoked by pressing C-c C-c once the document is formatted or via the respective entry in the Command menu.

AUCTEX will try to guess which type of viewer (DVI, PostScript or PDF) has to be used and what options are to be passed over to it. This decision is based on the output files present in the working directory as well as the class and style options used in the document. For example, if there is a DVI file in your working directory, a DVI viewer will be invoked. In case of a PDF file it will be a PDF viewer. If you specified a special paper format like ‘a5paper’ or use the ‘landscape’ option, this will be passed to the viewer by the appropriate options. Especially some DVI viewers depend on this kind of information in order to display your document correctly. In case you are using ‘pstricks’ or ‘psfrag’ in your document, a DVI viewer cannot display the contents correctly and a PostScript viewer will be invoked instead.

The information about which file types and style options are associated with which viewers and options for them is stored in the variables TeX-output-view-style and TeX-

view-style.

[Command] TeX-view

The command TeX-view, bound to C-c C-v, starts a viewer without confirmation. The viewer is started either on a region or the master file, depending on the last command issued. This is especially useful for jumping to the location corresponding to point in the DVI viewer when using TeX-source-specials-mode.

[User Option] TeX-output-view-style

List of output file extensions, style options and view options.

[User Option] TeX-view-style

List of style options and view options. This is the predecessor of TeX-output-view-style which does not allow the specification of output file extensions. It is used as a fallback in case none of the alternatives specified in TeX-output-view-style match. In case none of the entries in TeX-view-style match either, no suggestion for a viewer will be made.

7.2.2 Forward and inverse search

You can make use of forward and inverse searching if this is supported by your DVI viewer and you enabled TeX-source-specials-mode. See Section 7.1 [Commands], page 44, on how to do that. AUCTEX will automatically pass the necessary command line options to the viewer in order to display the page containing the content you are currently editing (forward search).

Upon opening the viewer you will be asked if you want to start a server process (Gnuserv or Emacs server) which is necessary for inverse search. This happens only if there is no server running already. You can customize the variable TeX-source-specials-view-start-server to inhibit the question and always or never start the server respectively.

[User Option] TeX-source-specials-view-start-server

If TeX-source-specials-mode is active and a DVI viewer is invoked, the default behavior is to ask if a server process should be started. Set this variable to t if the question should be inhibited and the server should always be started. Set it to nil if the server should never be started. Inverse search will not be available in the latter case.

Once the server and the viewer are running you can use a mouse click in the viewer to jump to the corresponding part of your document in Emacs (inverse search). Refer to the documentation of your viewer to find out what you have to do exactly. In xdvi you usually

have to use C-down-mouse-1.

For PDF output, forward search is availabe when using the pdfsync LaTEX package and xpdf as PDF viewer. With the pdfsync package forward search does not rely on source specials. Therefore you don’t have to bother about the provisions for source specials explained above. If document parsing is enabled, the functionality is usable immediately, e.g. by typing C-c C-v (TeX-view) which will open the viewer or bring it to front if it is already opened and display the output page corresponding to the position of point in the source file.

7.3 Catching the errors

Once you’ve formatted your document you may ‘debug’ it, i.e. browse through the errors (La)TEX reported.

[Command] TeX-next-error

(C-c `) Go to the next error reported by TEX. The view will be split in two, with the cursor placed as close as possible to the error in the top view. In the bottom view, the error message will be displayed along with some explanatory text.

Normally AUCTEX will only report real errors, but you may as well ask it to report ‘bad boxes’ and warnings as well.

[Command] TeX-toggle-debug-bad-boxes

(C-c C-t C-b) Toggle whether AUCTEX should stop at bad boxes (i.e. overfull and underfull boxes) as well as normal errors.

[Command] TeX-toggle-debug-warnings

(C-c C-t C-w) Toggle whether AUCTEX should stop at warnings as well as normal errors.

As default, AUCTEX will display that special ‘*help*’ buffer containing the error reported by TEX along with the documentation. There is however an ‘expert’ option, which allows you to display the real TEX output.

[User Option] TeX-display-help

When non-nil AUCTEX will automatically display a help text whenever an error is encountered using TeX-next-error (C-c `).

7.4 Checking for problems

Running TEX or LaTEX will only find regular errors in the document, not examples of bad style. Furthermore, description of the errors may often be confusing. The utility lacheck can be used to find style errors, such as forgetting to escape the space after an abbreviation or using ‘...’ instead of ‘\ldots’ and many other problems like that. You start lacheck with C-c C-c Check ^{⟨}\underline{\overline{RET}}^{⟩}. The result will be a list of errors in the ‘*compilation*’ buffer. You can go through the errors with C-x ` (next-error, see section “Compilation” in The Emacs Editor), which will move point to the location of the next error.

Another newer program which can be used to find errors is chktex. It is much more configurable than lacheck, but doesn’t find all the problems lacheck does, at least in its default configuration. You must install the programs before using them, and for chktex you may also need modify TeX-command-list unless you use its lacheck compatibility wrapper. You

can get lacheck from ‘<URL:ftp://ftp.ctan.org/tex-archive/support/lacheck/>’ or
alternatively chktex from ‘<URL:ftp://ftp.ctan.org/tex-archive/support/chktex/>’.

7.5 Controlling the output

A number of commands are available for controlling the output of an application running under AUCTEX

[Command] TeX-kill-job

(C-c C-k) Kill currently running external application. This may be either of TEX, LaTEX, previewer, BibTEX, etc.

[Command] TeX-recenter-output-buffer

(C-c C-l) Recenter the output buffer so that the bottom line is visible.

[Command] TeX-home-buffer

(C-c ^) Go to the ‘master’ file in the document associated with the current buffer, or if already there, to the file where the current process was started.

7.6 Cleaning intermediate and output files

[Command] TeX-clean

Remove generated intermediate files. In case a prefix argument is given, remove output files as well.

Canonical access to the function is provided by the ‘Clean’ and ‘Clean All’ entries in TeX-command-list, invokable with C-c C-c or the Command menu.

The patterns governing which files to remove can be adapted separately for each AUCTEX mode

by means of the variables plain-TeX-clean-intermediate-suffixes, plain-TeX-
clean-output-suffixes,  LaTeX-clean-intermediate-suffixes,
LaTeX-clean-output-suffixes, docTeX-clean-intermediate-suffixes, docTeX-
clean-output-suffixes, Texinfo-clean-intermediate-suffixes, Texinfo-
clean-output-suffixes, ConTeXt-clean-intermediate-suffixes and ConTeXt-
clean-output-suffixes.

[User Option] TeX-clean-confirm

Control if deletion of intermediate and output files has to be confirmed before it is actually done. If non-nil, ask before deleting files.

7.7 Documentation about macros and packages

[Command] TeX-doc

(C-c ?) Get documentation about macros, packages or TEX & Co. in general. The function will prompt for the name of a command or manual, providing a list of available keywords for completion. If point is on a command or word with available documentation, this will be suggested as default. The command can be invoked by the key binding mentioned above as well as the ‘Find Documentation...’ entry in the mode menu.

8 Multifile Documents

You may wish to spread a document over many files (as you are likely to do if there are multiple authors, or if you have not yet discovered the power of the outline commands (see Section 6.3 [Outline], page 43)). This can be done by having a “master” file in which you include the various files with the TEX macro ‘\input’ or the LaTEX macro ‘\include’. These files may also include other files themselves. However, to format the document you must run the commands on the top level master file.

When you, for example, ask AUCTEX to run a command on the master file, it has no way of knowing the name of the master file. By default, it will assume that the current file is the master file. If you insert the following in your ‘.emacs’ file AUCTEX will use a more advanced algorithm.

(setq-default TeX-master nil) ; Query for master file.

If AUCTEX finds the line indicating the end of the header in a master file (TeX-header-end), it can figure out for itself that this is a master file. Otherwise, it will ask for the name of the master file associated with the buffer. To avoid asking you again, AUCTEX will automatically insert the name of the master file as a file variable (see section “File Variables” in The Emacs Editor). You can also insert the file variable yourself, by putting the following text at the end of your files.

%%% Local Variables:
%%% TeX-master: "master"
%%% End:

You should always set this variable to the name of the top level document. If you always use the same name for your top level documents, you can set TeX-master in your ‘.emacs’ file.

(setq-default TeX-master "master") ; All master files called "master".

[User Option] TeX-master

The master file associated with the current buffer. If the file being edited is actually included from another file, then you can tell AUCTEX the name of the master file by setting this variable. If there are multiple levels of nesting, specify the top level file. If this variable is nil, AUCTEX will query you for the name. If the variable is t, then AUCTEX will assume the file is a master file itself. If the variable is shared, then AUCTEX will query for the name, but will not change the file.

[User Option] TeX-one-master

Regular expression matching ordinary TEX files. You should set this variable to match the name of all files, for which it is a good idea to append a TeX-master file variable entry automatically. When AUCTEX adds the name of the master file as a file variable, it does not need to ask next time you edit the file. If you dislike AUCTEX automatically modifying your files, you can set this variable to ‘"<none>"’. By default, AUCTEX will modify any file with an extension of ‘.tex’.

[Command] TeX-master-file-ask

(C-c _) Query for the name of a master file and add the respective File Variables (see section “File Variables” in The Emacs Editor) to the file for setting this variable permanently. AUCTEX will not ask for a master file when it encounters existing files. This function shall give you the possibility to insert the variable manually.

AUCTEX keeps track of macros, environments, labels, and style files that are used in a given document. For this to work with multifile documents, AUCTEX has to have a place to put the information about the files in the document. This is done by having an ‘auto’ subdirectory placed in the directory where your document is located. Each time you save a file, AUCTEX will write information about the file into the ‘auto’ directory. When you load a file, AUCTEX will read the information in the ‘auto’ directory about the file you loaded and the master file specified by TeX-master. Since the master file (perhaps indirectly) includes all other files in the document, AUCTEX will get information from all files in the document. This means that you will get from each file, for example, completion for all labels defined anywhere in the document.

AUCTEX will create the ‘auto’ directory automatically if TeX-auto-save is non-nil. Without it, the files in the document will not know anything about each other, except for the name of the master file. See Section 11.3 [Automatic Local], page 60.

[Command] TeX-save-document

(C-c C-d) Save all buffers known to belong to the current document.

[User Option] TeX-save-query

If non-nil, then query the user before saving each file with TeX-save-document.

9 Automatic Parsing of TEX Files

AUCTEX depends heavily on being able to extract information from the buffers by parsing them. Since parsing the buffer can be somewhat slow, the parsing is initially disabled. You are encouraged to enable them by adding the following lines to your ‘.emacs’ file.

     (setq TeX-parse-self t) ; Enable parse on load.
     (setq TeX-auto-save t) ; Enable parse on save.

The latter command will make AUCTEX store the parsed information in an ‘auto’ subdirectory in the directory each time the TEX files are stored, see Section 11.3 [Automatic Local], page 60. If AUCTEX finds the pre-parsed information when loading a file, it will not need to reparse the buffer. The information in the ‘auto’ directory is also useful for multifile documents, see Chapter 8 [Multifile], page 51, since it allows each file to access the parsed information from all the other files in the document. This is done by first reading the information from the master file, and then recursively the information from each file stored in the master file.

The variables can also be done on a per file basis, by changing the file local variables.

     %%% Local Variables:
     %%% TeX-parse-self: t
     %%% TeX-auto-save: t
     %%% End:

Even when you have disabled the automatic parsing, you can force the generation of style information by pressing C-c C-n. This is often the best choice, as you will be able to decide when it is necessary to reparse the file.

[User Option] TeX-parse-self

Parse file after loading it if no style hook is found for it.

[User Option] TeX-auto-save

Automatically save style information when saving the buffer.

[Command] TeX-normal-mode arg

(C-c C-n) Remove all information about this buffer, and apply the style hooks again. Save buffer first including style information. With optional argument, also reload the style hooks.

When AUCTEX saves your buffer, it can optionally convert all tabs in your buffer into spaces. Tabs confuse AUCTEX’s error message parsing and so should generally be avoided. However, tabs are significant in some environments, and so by default AUCTEX does not remove them. To convert tabs to spaces when saving a buffer, insert the following in your

‘.emacs’ file:
     (setq TeX-auto-untabify t)
[User Option] TeX-auto-untabify

Automatically remove all tabs from a file before saving it.

Instead of disabling the parsing entirely, you can also speed it significantly up by limiting the information it will search for (and store) when parsing the buffer. You can do this by setting the default values for the buffer local variables TeX-auto-regexp-list and TeX-

auto-parse-length in your ‘.emacs’ file.

;; Only parse LaTeX class and package information.
(setq-default TeX-auto-regexp-list 'LaTeX-auto-minimal-regexp-list)
;; The class and package information is usually near the beginning.
(setq-default TeX-auto-parse-length 2000)

This example will speed the parsing up significantly, but AUCTEX will no longer be able to provide completion for labels, macros, environments, or bibitems specified in the document, nor will it know what files belong to the document.

These variables can also be specified on a per file basis, by changing the file local variables.

%%% Local Variables:
%%% TeX-auto-regexp-list: TeX-auto-full-regexp-list
%%% TeX-auto-parse-length: 999999
%%% End:
[User Option] TeX-auto-regexp-list

List of regular expressions used for parsing the current file.

[User Option] TeX-auto-parse-length

Maximal length of TEX file that will be parsed.

The pre-specified lists of regexps are defined below. You can use these before loading AUCTEX by quoting them, as in the example above.

[Constant] TeX-auto-empty-regexp-list

Parse nothing

[Constant] LaTeX-auto-minimal-regexp-list

Only parse LaTEX class and packages.

[Constant] LaTeX-auto-label-regexp-list

Only parse LaTEX labels.

[Constant] LaTeX-auto-regexp-list

Parse common LaTEX commands.

[Constant] plain-TeX-auto-regexp-list

Parse common plain TEX commands.

[Constant] TeX-auto-full-regexp-list

Parse all TEX and LaTEX commands that AUCTEX can use.

10 Language Support

TEX and Emacs are usable for European (Latin, Cyrillic, Greek) based languages. Some LaTEX and EmacsLisp packages are available for easy typesetting and editing documents in European languages.

For CJK (Chinese, Japanese, and Korean) languages, Emacs or XEmacs with MULE (MULtilingual Enhancement to GNU Emacs) support is required. MULE is part of Emacs by default since Emacs 20. XEmacs has to be configured with the ‘--with-mule’ option. Special versions of TEX are needed for CJK languages: CTEX and ChinaTEX for Chinese, ASCII pTEX and NTT jTEX for Japanese, HLaTEX and kTEX for Korean. The CJK-LaTEX package is required for supporting multiple CJK scripts within a single document.

Note that Unicode is not fully supported in Emacs 21 and XEmacs 21. CJK characters are not usable. Please use the MULE-UCS EmacsLisp package or Emacs 22 (not released yet) if you need CJK.

10.1 Using AUCTEX with European Languages

10.1.1 Typing and Displaying Non-ASCII Characters

First you will need a way to write non-ASCII characters. You can either use macros, or teach TEX about the ISO character sets. I prefer the latter, it has the advantage that the usual standard emacs word movement and case change commands will work.

With LaTEX2e, just add ‘\usepackage[latin1]{inputenc}’. Other languages than Western European ones will probably have other encoding needs.

To be able to display non-ASCII characters you will need an appropriate font and a version of GNU Emacs capable of displaying 8-bit characters (e.g. Emacs 21). The manner in which this is supported differs between Emacsen, so you need to take a look at your respective documentation.

A compromise is to use an European character set when editing the file, and convert to TEX macros when reading and writing the files.

‘iso-cvt.el’

Much like ‘iso-tex.el’ but is bundled with Emacs 19.23 and later.

‘x-compose.el’

Similar package bundled with new versions of XEmacs.

‘X-Symbol’

a much more complete package for both Emacs and XEmacs that can also handle a lot of mathematical characters and input methods.

10.1.2 Style Files for Different Languages

AUCTEX supports style files for several languages. Each style file may modify AUCTEX to better support the language, and will run a language specific hook that will allow you to for example change ispell dictionary, or run code to change the keyboard remapping. The following will for example choose a Danish dictionary for documents including ‘\usepackage[danish]{babel}’. This requires parsing to be enabled, see Chapter 9 [Parsing Files], page 53.

     (add-hook 'TeX-language-dk-hook
       (lambda () (ispell-change-dictionary "danish")))

The following style files are recognized:

‘czech’ Runs style hook TeX-language-cz-hook. Pressing ^{⟨}\underline{\overline{"}}^{⟩} will insert ‘\uv\{’ and ‘\}’ depending on context.

‘danish’ Runs style hook TeX-language-dk-hook. Pressing ^{⟨}\underline{\overline{"}}^{⟩} will insert ‘"`’ and ‘"'’ depending on context. Typing ^{⟨}\underline{\overline{-}}^{⟩} twice will insert ‘"=’, i.e. a hyphen string allowing hyphenation in the composing words.

‘dutch’  Runs style hook TeX-language-nl-hook.
‘german’

‘ngerman’ Runs style hook TeX-language-de-hook. Gives ‘"’ word syntax, makes the ^{⟨}\underline{\overline{"}}^{⟩} key insert a literal ‘"’. Pressing the key twice will give you opening or closing German quotes (‘"`’ or ‘"'’). Typing ^{⟨}\underline{\overline{-}}^{⟩} twice will insert ‘"=’, three times ‘--’.

‘frenchb’
‘francais’

Runs style hook TeX-language-fr-hook. Pressing ^{⟨}\underline{\overline{"}}^{⟩} will insert ‘\\og’ and ‘\\fg’ depending on context. Note that the language name for customizing

          TeX-quote-language-alist is ‘french’.

‘italian’ Runs style hook TeX-language-it-hook. Pressing ^{⟨}\underline{\overline{"}}^{⟩} will insert ‘"<’ and ‘">’ depending on context.

‘polish’ Runs style hook TeX-language-pl-hook. Gives ‘"’ word syntax and makes the ⟨\underline{\overline{"}}^{⟩} key insert a literal ‘"’. Pressing ^{⟨}\underline{\overline{"}}^{⟩} twice will insert ‘"`’ or ‘"'’ depending on context.

‘polski’ Runs style hook TeX-language-pl-hook. Makes the ^{⟨}\underline{\overline{"}}^{⟩} key insert a literal ‘"’. Pressing ^{⟨}\underline{\overline{"}}^{⟩} twice will insert ‘,,’ or ‘''’ depending on context.

‘slovak’ Runs style hook TeX-language-sk-hook. Pressing ^{⟨}\underline{\overline{"}}^{⟩} will insert ‘\uv\{’ and ‘\}’ depending on context.

‘swedish’ Runs style hook TeX-language-sv-hook. Pressing ^{⟨}\underline{\overline{"}}^{⟩} will insert ‘''’. Typing ⟨\underline{\overline{-}}^{⟩} twice will insert ‘"=’, three times ‘--’.

Replacement of language-specific hyphen strings like ‘"=’ with dashes does not require to type ^{⟨}\underline{\overline{-}}^{⟩} three times in a row. You can put point after the hypen string anytime and trigger the replacement by typing ^{⟨}\underline{\overline{-}}^{⟩}.

In case you are not satisfied with the suggested behavior of quote and hyphen insertion you can change it by customizing the variables TeX-quote-language-alist and LaTeX-

babel-hyphen-language-alist respectively.
[User Option] TeX-quote-language-alist

Used for overriding the default language-specific quote insertion behavior. This is an alist where each element is a list consisting of four items. The first item is the name of the language in concern as a string. See the list of supported languages above. The second item is the opening quotation mark. The third item is the closing quotation mark. Opening and closing quotation marks can be specified directly as strings or as functions returning a string. The fourth item is a boolean controlling quote insertion. It should be non-nil if if the special quotes should only be used after inserting a literal ‘"’ character first, i.e. on second key press.

[User Option] LaTeX-babel-hyphen-language-alist

Used for overriding the behavior of hyphen insertion for specific languages. Every element in this alist is a list of three items. The first item should specify the affected language as a string. The second item denotes the hyphen string to be used as a string. The third item, a boolean, controls the behavior of hyphen insertion and should be non-nil if the special hyphen should be inserted after inserting a literal ‘-’ character, i.e. on second key press.

The defaults of hyphen insertion are defined by the variables LaTeX-babel-hyphen and

LaTeX-babel-hyphen-after-hyphen respectively.
[User Option] LaTeX-babel-hyphen

String to be used when typing ^{⟨}\underline{\overline{-}}^{⟩}. This usually is a hyphen alternative or hyphenation aid provided by ‘babel’ and the related language style files, like ‘"=’, ‘"~’ or ‘"-’. Set it to an empty string or nil in order to disable language-specific hyphen insertion.

[User Option] LaTeX-babel-hyphen-after-hyphen

Control insertion of hyphen strings. If non-nil insert normal hyphen on first key press and swap it with the language-specific hyphen string specified in the variable LaTeX-babel-hyphen on second key press. If nil do it the other way round.

10.2 Using AUCTEX with Japanese TEX

To write Japanese text with AUCTEX, you need to have versions of TEX and Emacs that support Japanese. There exist at least two variants of TEX for Japanese text (NTT jTEX and ASCII pTEX). AUCTEX can be used with MULE (MULtilingual Enhancement to GNU Emacs) supported Emacsen.

To use the Japanese TEX variants, simply activate japanese-plain-tex-mode or japanese-latex-mode and everything should work. If not, send mail to Masayuki Ataka ‘<ataka@milk.freemail.ne.jp>’, who kindly donated the code for supporting Japanese in AUCTEX. None of the primary AUCTEX maintainers understand Japanese, so they cannot help you.

If you usually use AUCTEX in Japanese, setting the following variables is useful.

[User Option] TeX-default-mode

Mode to enter for a new file when it cannott be determined whether the file is plain TEX or LaTEX or what. If you want to enter Japanese LaTEX mode whenever this may happen, set the variable like this:

     (setq TeX-default-mode 'japanese-latex-mode)
[User Option] japanese-TeX-command-default

The default command for TeX-command in Japanese TEX mode. The default value is ‘"pTeX"’.

[User Option] japanese-LaTeX-command-default

The default command for TeX-command in Japanese LaTEX mode. The default value is ‘"LaTeX"’.

[User Option] japanese-LaTeX-default-style

The default style/class when creating a new Japanese LaTEX document. The default value is ‘"jarticle"’.

See ‘tex-jp.el’ for more information.

11 Automatic Customization

Since AUCTEX is so highly customizable, it makes sense that it is able to customize itself. The automatic customization consists of scanning TEX files and extracting symbols, environments, and things like that.

The automatic customization is done on three different levels. The global level is the level shared by all users at your site, and consists of scanning the standard TEX style files, and any extra styles added locally for all users on the site. The private level deals with those style files you have written for your own use, and use in different documents. You may have a ‘~/lib/TeX/’ directory where you store useful style files for your own use. The local level is for a specific directory, and deals with writing customization for the files for your normal TEX documents.

If compared with the environment variable TEXINPUTS, the global level corresponds to the directories built into TEX. The private level corresponds to the directories you add yourself, except for ‘.’, which is the local level.

By default AUCTEX will search for customization files in all the global, private, and local style directories, but you can also set the path directly. This is useful if you for example want to add another person’s style hooks to your path. Please note that all matching files found in TeX-style-path are loaded, and all hooks defined in the files will be executed.

[User Option] TeX-style-path

List of directories to search for AUCTEX style files. Each must end with a slash.

By default, when AUCTEX searches a directory for files, it will recursively search through subdirectories.

[User Option] TeX-file-recurse

Whether to search TEX directories recursively: nil means do not recurse, a positive integer means go that far deep in the directory hierarchy, t means recurse indefinitely.

By default, AUCTEX will ignore files name ‘.’, ‘..’, ‘SCCS’, ‘RCS’, and ‘CVS’.

[User Option] TeX-ignore-file

Regular expression matching file names to ignore.

These files or directories will not be considered when searching for TEX files in a directory.

11.1 Automatic Customization for the Site

Assuming that the automatic customization at the global level was done when AUCTEX was installed, your choice is now: will you use it? If you use it, you will benefit by having access to all the symbols and environments available for completion purposes. The drawback is slower load time when you edit a new file and perhaps too many confusing symbols when you try to do a completion.

You can disable the automatic generated global style hooks by setting the variable TeX-

auto-global to nil.
[User Option] TeX-macro-global

Directories containing the site’s TEX style files.

[User Option] TeX-style-global

Directory containing hand generated TEX information. Must end with a slash.

These correspond to TEX macros shared by all users of a site.

[User Option] TeX-auto-global

Directory containing automatically generated information.

For storing automatic extracted information about the TEX macros shared by all users of a site.

11.2 Automatic Customization for a User

You should specify where you store your private TEX macros, so AUCTEX can extract their information. The extracted information will go to the directories listed in TeX-auto-

private

Use M-x TeX-auto-generate to extract the information.

[User Option] TeX-macro-private

Directories where you store your personal TEX macros. Each must end with a slash.

This defaults to the directories listed in the ‘TEXINPUTS’ and ‘BIBINPUTS’ environment variables.

[User Option] TeX-auto-private

List of directories containing automatically generated information. Must end with a slash.

These correspond to the personal TEX macros.

[Command] TeX-auto-generate TEX AUTO

(M-x TeX-auto-generate) Generate style hook for TEX and store it in AUTO. If TEX is a directory, generate style hooks for all files in the directory.

[User Option] TeX-style-private

List of directories containing hand generated information. Must end with a slash.

These correspond to the personal TEX macros.

11.3 Automatic Customization for a Directory

AUCTEX can update the style information about a file each time you save it, and it will do this if the directory TeX-auto-local exist. TeX-auto-local is by default set to ‘"auto/"’, so simply creating an ‘auto’ directory will enable automatic saving of style information.

The advantage of doing this is that macros, labels, etc. defined in any file in a multifile document will be known in all the files in the document. The disadvantage is that saving will be slower. To disable, set TeX-auto-local to nil.

[User Option] TeX-style-local

Directory containing hand generated TEX information. Must end with a slash.

These correspond to TEX macros found in the current directory.

[User Option] TeX-auto-local

Directory containing automatically generated TEX information. Must end with a slash. These correspond to TEX macros found in the current directory.

12 Writing Your own Style Support

See Chapter 11 [Automatic], page 59, for a discussion about automatically generated global, private, and local style files. The hand generated style files are equivalent, except that they by default are found in ‘style’ directories instead of ‘auto’ directories.

If you write some useful support for a public TEX style file, please send it to us.

12.1 A Simple Style File

Here is a simple example of a style file.

     ;;; book.el - Special code for book style.
     (TeX-add-style-hook
      "book"
      (lambda () (setq LaTeX-largest-level
       (LaTeX-section-level ("chapter")))))

This file specifies that the largest kind of section in a LaTEX document using the book document style is chapter. The interesting thing to notice is that the style file defines an (anonymous) function, and adds it to the list of loaded style hooks by calling TeX-add-

style-hook.

The first time the user indirectly tries to access some style specific information, such as the largest sectioning command available, the style hooks for all files directly or indirectly read by the current document is executed. The actual files will only be evaluated once, but the hooks will be called for each buffer using the style file.

[Function] TeX-add-style-hook style hook

Add hook to the list of functions to run when we use the TEX file style.

12.2 Adding Support for Macros

The most common thing to define in a style hook is new symbols (TEX macros). Most likely along with a description of the arguments to the function, since the symbol itself can be defined automatically.

Here are a few examples from ‘latex.el’.

     (TeX-add-style-hook
      "latex"
      (lambda ()
        (TeX-add-symbols
         '("arabic" TeX-arg-counter)
         '("label" TeX-arg-define-label)
         '("ref" TeX-arg-label)
         '("newcommand" TeX-arg-define-macro [ "Number of arguments" ] t)
         '("newtheorem" TeX-arg-define-environment
           [ TeX-arg-environment "Numbered like" ]
           t [ TeX-arg-counter "Within counter" ]))))
[Function] TeX-add-symbols symbol . . .

Add each symbol to the list of known symbols.

Each argument to TeX-add-symbols is a list describing one symbol. The head of the list is the name of the symbol, the remaining elements describe each argument.

If there are no additional elements, the symbol will be inserted with point inside braces. Otherwise, each argument of this function should match an argument of the TEX macro. What is done depends on the argument type.

If a macro is defined multiple times, AUCTEX will chose the one with the longest definition (i.e. the one with the most arguments).

Thus, to overwrite

     '("tref" 1) ; one argument

you can specify

     '("tref" TeX-arg-label ignore) ; two arguments

ignore is a function that does not do anything, so when you insert a ‘tref’ you will be prompted for a label and no more.

string Use the string as a prompt to prompt for the argument.

number Insert that many braces, leave point inside the first.

nil Insert empty braces.

t Insert empty braces, leave point between the braces.

other symbols

Call the symbol as a function. You can define your own hook, or use one of the predefined argument hooks.

list If the car is a string, insert it as a prompt and the next element as initial input. Otherwise, call the car of the list with the remaining elements as arguments.

vector Optional argument. If it has more than one element, parse it as a list, otherwise parse the only element as above. Use square brackets instead of curly braces, and is not inserted on empty user input.

A lot of argument hooks have already been defined. The first argument to all hooks is a flag indicating if it is an optional argument. It is up to the hook to determine what to do with the remaining arguments, if any. Typically the next argument is used to overwrite the default prompt.

TeX-arg-conditional

Implements if EXPR THEN ELSE. If EXPR evaluates to true, parse THEN as an argument list, else parse ELSE as an argument list.

TeX-arg-literal

Insert its arguments into the buffer. Used for specifying extra syntax for a macro.

TeX-arg-free

Parse its arguments but use no braces when they are inserted.

TeX-arg-eval

Evaluate arguments and insert the result in the buffer.

TeX-arg-label

Prompt for a label completing with known labels.

TeX-arg-macro

Prompt for a TEX macro with completion.

TeX-arg-environment

Prompt for a LaTEX environment with completion.

TeX-arg-cite

Prompt for a BibTEX citation.

TeX-arg-counter

Prompt for a LaTEX counter.

TeX-arg-savebox

Prompt for a LaTEX savebox.

TeX-arg-file

Prompt for a filename in the current directory, and use it without the extension.

TeX-arg-input-file

Prompt for the name of an input file in TEX’s search path, and use it without the extension. Run the style hooks for the file.

TeX-arg-define-label

Prompt for a label completing with known labels. Add label to list of defined labels.

TeX-arg-define-macro

Prompt for a TEX macro with completion. Add macro to list of defined macros.

TeX-arg-define-environment

Prompt for a LaTEX environment with completion. Add environment to list of defined environments.

TeX-arg-define-cite

Prompt for a BibTEX citation.

TeX-arg-define-counter

Prompt for a LaTEX counter.

TeX-arg-define-savebox

Prompt for a LaTEX savebox.

TeX-arg-corner

Prompt for a LaTEX side or corner position with completion.

TeX-arg-lr

Prompt for a LaTEX side with completion.

TeX-arg-tb

Prompt for a LaTEX side with completion.

TeX-arg-pagestyle

Prompt for a LaTEX pagestyle with completion.

TeX-arg-verb

Prompt for delimiter and text.

TeX-arg-pair

Insert a pair of numbers, use arguments for prompt. The numbers are surrounded by parentheses and separated with a comma.

TeX-arg-size

Insert width and height as a pair. No arguments.

TeX-arg-coordinate

Insert x and y coordinates as a pair. No arguments.

If you add new hooks, you can assume that point is placed directly after the previous argument, or after the macro name if this is the first argument. Please leave point located after the argument you are inserting. If you want point to be located somewhere else after all hooks have been processed, set the value of exit-mark. It will point nowhere, until the argument hook sets it.

12.3 Adding Support for Environments

Adding support for environments is very much like adding support for TEX macros, except that each environment normally only takes one argument, an environment hook. The example is again a short version of ‘latex.el’.

     (TeX-add-style-hook
      "latex"
      (lambda ()
        (LaTeX-add-environments
         '("document" LaTeX-env-document)
         '("enumerate" LaTeX-env-item)
         '("itemize" LaTeX-env-item)
         '("list" LaTeX-env-list))))

The only hook that is generally useful is LaTeX-env-item, which is used for environments that contain items. It is completely up to the environment hook to insert the environment, but the function LaTeX-insert-environment may be of some help. The hook will be called with the name of the environment as its first argument, and extra arguments can be provided by adding them to a list after the hook.

For simple environments with arguments, for example defined with ‘\newenvironment’, you can make AUCTEX prompt for the arguments by giving the prompt strings in the call to LaTeX-add-environments. For example, if you have defined a loop environment with the three arguments from, to, and step, you can add support for them in a style file.

     %% loop.sty
     \newenvironment{loop}[3]{...}{...}
     ;; loop.el
     (TeX-add-style-hook
      "loop"

      (lambda ()
        (LaTeX-add-environments
         '("loop" "From" "To" "Step"))))

If an environment is defined multiple times, AUCTEX will chose the one with the longest definition. Thus, if you have an enumerate style file, and want it to replace the standard LaTEX enumerate hook above, you could define an ‘enumerate.el’ file as follows, and place it in the appropriate style directory.

     (TeX-add-style-hook
      "latex"
      (lambda ()
        (LaTeX-add-environments
         '("enumerate" LaTeX-env-enumerate foo))))
     (defun LaTeX-env-enumerate (environment &optional ignore) ...)

The symbol foo will be passed to LaTeX-env-enumerate as the second argument, but since we only added it to overwrite the definition in ‘latex.el’ it is just ignored.

[Function] LaTeX-add-environments env . . .

Add each env to list of loaded environments.

[Function] LaTeX-insert-environment env [ extra ]

Insert environment of type env, with optional argument extra.

12.4 Adding Other Information

You can also specify bibliographical databases and labels in the style file. This is probably of little use, since this information will usually be automatically generated from the TEX file anyway.

[Function] LaTeX-add-bibliographies bibliography . . .

Add each bibliography to list of loaded bibliographies.

[Function] LaTeX-add-labels label . . .

Add each label to the list of known labels.

12.5 Automatic Extraction of New Things

The automatic TEX information extractor works by searching for regular expressions in the TEX files, and storing the matched information. You can add support for new constructs to the parser, something that is needed when you add new commands to define symbols.

For example, in the file ‘macro.tex’ I define the following macro.

     \newcommand{\newmacro}[5]{%
     \def#1{#3\index{#4@#5~cite{#4}}\nocite{#4}}%
     \def#2{#5\index{#4@#5~cite{#4}}\nocite{#4}}%
     }

AUCTEX will automatically figure out that ‘newmacro’ is a macro that takes five arguments. However, it is not smart enough to automatically see that each time we use the macro, two new macros are defined. We can specify this information in a style hook file.

;;; macro.el --- Special code for my own macro file.
;;; Code:
(defvar TeX-newmacro-regexp
  '("\\\\newmacro{\\\\\\([a-zA-Z]+\\)}{\\\\\\([a-zA-Z]+\\)}"
    (1 2) TeX-auto-multi)
  "Matches \newmacro definitions.")
(defvar TeX-auto-multi nil
  "Temporary for parsing \\newmacro definitions.")
(defun TeX-macro-cleanup ()
  "Move symbols from `TeX-auto-multi' to `TeX-auto-symbol'."
  (mapcar (lambda (list)
    (mapcar (lambda (symbol)
      (setq TeX-auto-symbol
    (cons symbol TeX-auto-symbol)))
    list))
  TeX-auto-multi))
(defun TeX-macro-prepare ()
  "Clear `Tex-auto-multi' before use."
  (setq TeX-auto-multi nil))
(add-hook 'TeX-auto-prepare-hook 'TeX-macro-prepare)
(add-hook 'TeX-auto-cleanup-hook 'TeX-macro-cleanup)
(TeX-add-style-hook
 "macro"
 (lambda ()
   (TeX-auto-add-regexp TeX-newmacro-regexp)
   (TeX-add-symbols '("newmacro"
      TeX-arg-macro
      (TeX-arg-macro "Capitalized macro: \\")
      t
      "BibTeX entry: "
      nil))))
;;; macro.el ends here

When this file is first loaded, it adds a new entry to TeX-newmacro-regexp, and defines a function to be called before the parsing starts, and one to be called after the parsing is done. It also declares a variable to contain the data collected during parsing. Finally, it adds a style hook which describes the ‘newmacro’ macro, as we have seen it before.

So the general strategy is: Add a new entry to TeX-newmacro-regexp. Declare a variable to contain intermediate data during parsing. Add hook to be called before and after parsing.

In this case, the hook before parsing just initializes the variable, and the hook after parsing collects the data from the variable, and adds them to the list of symbols found.

[Variable] TeX-auto-regexp-list

List of regular expressions matching TEX macro definitions. The list has the following format ((REGEXP MATCH TABLE) …), that is, each entry is a list with three elements. REGEXP. Regular expression matching the macro we want to parse. MATCH. A number or list of numbers, each representing one parenthesized subexpression matched by REGEXP. TABLE. The symbol table to store the data. This can be a function, in which case the function is called with the argument MATCH. Use TeX-match-buffer to get match data. If it is not a function, it is presumed to be the name of a variable containing a list of match data. The matched data (a string if MATCH is a number, a list of strings if MATCH is a list of numbers) is put in front of the table.

[Variable] TeX-auto-prepare-hook nil

List of functions to be called before parsing a TEX file.

[Variable] TeX-auto-cleanup-hook nil

List of functions to be called after parsing a TEX file.

Appendix A Copying this Manual

The full license text can be read here:

A.1 GNU Free Documentation License

Version 1.2, November 2002

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000,2001,2002 Free Software Foundation, Inc. 51 Franklin St, Fifth Floor, Boston, MA 02110-1301, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none. The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words. A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”. Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only. The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text. A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition. The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING

You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3.

You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY

If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects.

If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages.

If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public.

It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS

You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version:

A. Use in the Title Page (and on the covers, if any) a title distinct from that of the Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided for under this License. Any other attempt to copy, modify, sublicense or distribute the Document is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.2
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with...Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Appendix B Changes and New Features

News in 11.84

• There have been problems with the ‘-without-texmf-dir’ option to ‘configure’ when the value of ‘-with-kpathsea-sep’ was set or determined for an installation system with a default different from that of the runtime system. with-kpathsea-sep has been removed; the setting is now usually determined at runtime. Due to this and other problems, preview-latex in the released XEmacs package failed under Windows or with anything except recent 21.5 XEmacsen.

• AUCTEX and preview-latex have been changed in order to accommodate file names containing spaces. preview-latex now tolerates bad PostScript code polluting the stack (like some Omega fonts).

• ‘preview.sty’ had in some cases failed to emit PostScript header specials.

• Support for folding of comments was added.

• The polish language option of the babel LaTeX package as well as the polski LaTeX package are now supported. Most notably this means that AUCTeX will help to insert quotation marks as defined by polish.sty ("`..."') and polski.sty (,,...'').

• The TEX tool bar is now available and enabled by default in plain TEX mode. See Chapter 7 [Running TeX and friends], page 44.

• Bug fix in the display of math subscripts and superscripts.

• Bug fix TeX-doc for Emacs 21.

• There has been quite a number of other bug fixes to various features and documentation across the board.

News in 11.83

• The new function TeX-doc provides easy access to documentation about commands and packages or information related to TEX and friends in general. See Section 7.7 [Documentation], page 50.

• You can now get rid of generated intermediate and output files by means of the new

‘Clean’ and ‘Clean All’ entries in TeX-command-list accessible with C-c C-c or the

Command menu. See Section 7.6 [Cleaning], page 49.

• Support for forward search with PDF files was added. That means you can jump to a place in the output file corresponding to the position in the source file. See Section 7.2 [Viewing], page 46. Adding support for this feature required the default value of the variable TeX-output-view-style to be changed. Please make sure you either remove any customizations overriding the new default or incorporate the changes into your customizations if you want to use this feature.

• TEX error messages of the -file-line-error kind are now understood in AUCTEX and preview-latex (parsers are still separate).

• Bug fix in XyMTEX support.

• The LaTEX tool bar is now enabled by default. See Chapter 7 [Running TeX and friends], page 44.

News in 11.82

• Support for the MinionPro LaTeX package was added.

• Warnings and underfull/overfull boxes are now being indicated in the echo area after a LaTEX run, if the respective debugging options are activated with TeX-toggle-debug-

warnings (C-c C-t C-w) or TeX-toggle-debug-bad-boxes (C-c C-t C-b). In this case

TeX-next-error will find these warnings in addition to normal errors. The key binding C-c C-w for TeX-toggle-debug-bad-boxes (which was renamed from

TeX-toggle-debug-boxes) now is deprecated.

• AUCTEX now can automatically insert a pair of braces after typing ^{⟨} ^{⟩} or ^{⟨}\underline{\overline{^}}^{⟩} in math constructs if the new variable TeX-electric-sub-and-superscript is set to a non-nil value.

• Some language-specific support for French was added. There now is completion support for the commands provided by the ‘frenchb’ (and ‘francais’) options of the babel LaTEX package and easier input of French quotation marks (\\og ...\\fg) which can now be inserted by typing ^{⟨}\underline{\overline{"}}^{⟩}.

• Completion support for options of some LaTeX packages was added.

• Already in version 11.81 the way to activate AUCTEX changed substantially. This should now be done with (load "auctex.el" nil t t) instead of the former (require 'tex-site). Related to this change ‘tex-mik.el’ does not load ‘tex-site.el’ anymore. That means if you used only (require 'tex-mik) in order to activate AUCTEX, you have to add (load "auctex.el" nil t t) before the latter statement. More detailed information can be found in the installation instructions.

• Handling of verbatim constructs was consolidated across AUCTeX. This resulted

in the font-latex-specific variables font-latex-verb-like-commands, font-latex-
verbatim-macros, and font-latex-verbatim-environments being removed and
the more general variables LaTeX-verbatim-macros-with-delims, LaTeX-verbatim-
macros-with-braces, and LaTeX-verbatim-environments being added.

• The output of a BibTEX run is now checked for warnings and errors, which are reported in the echo area.

• The aliases for font-latex-title-fontify were removed. Use font-latex-fontify-
sectioning instead.

• The problem that Japanese macros where broken across lines was fixed.

• Various bug fixes.

News in 11.81

• LaTeX-mark-section now marks subsections of a given section as well. The former behavior is available via the prefix argument.

• preview-latex which was previously available separately became a subsystem of AUCTEX. There is no documented provision for building or installing preview-latex separately. It is still possible to use and install AUCTEX without preview-latex, however.

• The installation procedures have been overhauled and now also install startup files as part of the process (those had to be copied manually previously). You are advised to remove previous installations of AUCTEX and preview-latex before starting the installation procedure. A standard installation from an unmodified tarball no longer requires Makeinfo or Perl. Also note that the way AUCTEX is supposed to be activated changed. Instead of

(require 'tex-site) you should now use (load "auctex.el" nil t t). While the

former method may still work, the new method has the advantage that you can deactivate a preactivated AUCTEX with the statement (unload-feature 'tex-site) before any of its modes have been used. This may be important especially for site-wide installations.

• Support for the babel LaTEX package was added.

• Folding a buffer now ensures that the whole buffer is fontified before the actual folding is carried out. If this results in unbearably long execution times, you can fall back to the old behavior of relying on stealth font locking to do this job in the background by customizing the variable TeX-fold-force-fontify.

• Folded content now reveals part of its original text in a tooltip or the echo area when hovering with the mouse pointer over it.

• The language-specific insertion of quotation marks was generalized. The

variables LaTeX-german-open-quote, LaTeX-german-close-quote, LaTeX-german-
quote-after-quote, LaTeX-italian-open-quote, LaTeX-italian-close-quote,

and LaTeX-italian-quote-after-quote are now obsolete. If you are not satisfied with the default settings, you should customize TeX-quote-language-alist instead.

• Similar to language-specific quote insertion, AUCTEX now helps you with hyphens in different languages as well. See Section 10.1 [European], page 55, for details.

• Fill problems in Japanese text introduced in AUCTEX 11.55 were fixed. AUCTEX tries not to break lines between 1-byte and 2-byte chars. These features will work in Chinese text, too.

• The scaling factor of the fontification of sectioning commands can now be customized using the variable font-latex-fontify-sectioning. This variable was previously called font-latex-title-fontify; In this release we provide an alias but this will disappear in one of the the next releases. The faces for the sectioning commands are now

called font-latex-sectioning-N-face (N=0. . .5) instead of font-latex-title-N-

face (N=1…4). Analogously the names of the variables holding the related keyword

lists were changed from font-latex-title-N-keywords to font-latex-sectioning-

N-keywords. See Section 6.1 [Font Locking], page 35, for details. Make sure to adjust your customizations.

• Titles in beamer slides marked by the “\frametitle” command are know displayed with the new face font-latex-slide-title-face. You can add macros to be highlighted

with this face to font-latex-match-slide-title-keywords.

• Of course a lot of bugs have been fixed.

News in 11.55

• A bug was fixed which lead to the insertion of trailing whitespace during filling. In particular extra spaces were added to sentence endings at the end of lines. You can make this whitespace visible by setting the variable show-trailing-whitespace to t.

If you want to delete all trailing whitespace in a buffer, type M-x delete-trailing-

whitespace RET.

• A bug was fixed which lead to a ‘*Compile-Log*’ buffer popping up when the first LaTEX file was loaded in an Emacs session.

• On some systems the presence of an outdated Emacspeak package lead to the

error message ‘File mode specification error: (error "Variable binding depth

exceeds max-specpdl-size")’. Precautions were added which prevent this error from happening. But nevertheless, it is advised to upgrade or uninstall the outdated Emacspeak package.

• The value of TeX-macro-global is not determined during configuration anymore but at load time of AUCTEX. Consequently the associated configuration option

‘--with-tex-input-dirs’ was removed.

• Support for the LaTEX Japanese classes ‘jsarticle’ and ‘jsbook’ was added.

News in 11.54

• The parser (used e.g. for TeX-auto-generate-global) was extended to recognize keywords common in LaTEX packages and classes, like “\DeclareRobustCommand” or “\RequirePackage”. Additionally a bug was fixed which led to duplicate entries in AUCTEX style files.

• Folding can now be done for paragraphs and regions besides single constructs and the whole buffer. With the new TeX-fold-dwim command content can both be hidden and shown with a single key binding. In course of these changes new key bindings for unfolding commands where introduced. The old bindings are still present but will be phased out in future releases.

• Info files of the manual now have a .info extension.

• There is an experimental tool bar support now. It is not activated by default. If you

want to use it, add

     (add-hook 'LaTeX-mode-hook 'LaTeX-install-toolbar)

to your init file.

• The manual now contains a new chapter “Quick Start”. It explains the main features and how to use them, and should be enough for a new user to start using AUCTEX.

• A new section “Font Locking” was added to the manual which explains syntax highlighting in AUCTEX and its customization. Together with the sections related to folding and outlining, the section is part of the new chapter “Display”.

• Keywords for syntax highlighting of LaTEX constructs to be typeset in bold, italic or typewriter fonts may now be customized. Besides the builtin classes, new keyword classes may be added by customizing the variable ‘font-latex-user-keyword-classes’. The customization options can be found in the customization group ‘font-latex-keywords’.

• Verbatim content is now displayed with the ‘fixed-pitch’ face. (GNU Emacs only)

• Syntax highlighting should not spill out of verbatim content anymore. (GNU Emacs only)

• Verbatim commands like ‘\verb|...|’ will not be broken anymore during filling.

• You can customize the completion for graphic files with LaTeX-includegraphics-

  read-file.

• Support for the LaTEX packages ‘url’, ‘listings’, ‘jurabib’ and ‘csquotes’ was added with regard to command completion and syntax highlighting.

• Performance of fontification and filling was improved.

• Insertion of nodes in Texinfo mode now supports completion of existing node names.

• Setting the variable LaTeX-float to nil now means that you will not be prompted for the float position of figures and tables. You can get the old behaviour of nil by setting the variable to "", i.e. an empty string. See also Section 4.4.2 [Floats], page 26.

• The XEmacs-specific bug concerning overlays-at was fixed.

• Lots of bug fixes.

News in 11.53

• The LaTEX math menu can include Unicode characters if your Emacs built supports it. See the variable LaTeX-math-menu-unicode, Section 5.1 [Mathematics], page 28.

• Bug fixes for XEmacs.

• Completion for graphic files in the TeX search path has been added.

• start is used for the viewer for MiKTEX and fpTEX.

• The variable TeX-fold-preserve-comments can now be customized to deactivate folding in comments.

News in 11.52

• Installation and menus under XEmacs work again (maybe for the first time).

• Fontification of subscripts and superscripts is now disabled when the fontification engine is not able to support it properly.

• Bug fixes in the build process.

News in 11.51

• PDFTeX and Source Special support did not work with ConTeXt, this has been fixed. Similar for Source Special support under Windows.

• Omega support has been added.

• Bug fixes in the build process.

• TeX-fold now supports folding of environments in Texinfo mode.

News in 11.50

• The use of source specials when processing or viewing the document can now be controlled with the new TeX-source-specials minor mode which can be toggled via an entry in the Command menu or the key binding C-c C-t C-s. If you have customized the variable TeX-command-list, you have to re-initialize it for this to work. This means to open a customization buffer for the variable by typing M-x customize-variable RET TeX-command-list RET, selecting “Erase Customization” and do your customization again with the new default.

• The content of the command menu now depends on the mode (plain TEX, LaTEX, ConTEXt etc.). Any former customization of the variable TeX-command-list has to be erased. Otherwise the command menu and the customization will not work correctly.

• Support for hiding and auto-revealing macros, e.g. footnotes or citations, and environments in a buffer was added, Section 6.2 [Folding], page 40.

• You can now control if indentation is done upon typing^{⟨}\underline{\overline{RET}}^{⟩} by customizing the variable TeX-newline-function, Section 5.4 [Indenting], page 31.

• Limited support for doc.sty and ltxdoc.cls (‘dtx’ files) was added. The new docTEX mode provides functionality for editing documentation parts. This includes formatting (indenting and filling), adding and completion of macros and environments while staying in comments as well as syntax highlighting. (Please note that the mode is not finished yet. For example syntax highlighting does not work yet in XEmacs.)

• For macro completion in docTEX mode the AUCTEX style files ‘doc.el’, ‘ltxdoc.el’ and ‘ltx-base.el’ were included. The latter provides general support for low-level LaTEX macros and may be used with LaTEX class and style files as well. It is currently not loaded automatically for those files.

• Support for ConTEXt with a separate ConTEXt mode is now included. Macro definitions for completion are available in Dutch and English.

• The filling and indentation code was overhauled and is now able to format commented parts of the source syntactically correct. Newly available functionality and customization options are explained in the manual.

• Filling and indentation in XEmacs with preview-latex and activated previews lead to the insertion of whitespace before multi-line previews. AUCTEX now contains facilities to prevent this problem.

• If TeX-master is set to t, AUCTEX will now query for a master file only when a new file is opened. Existing files will be left alone. The new function TeX-master-file-ask (bound to C-c _is provided for adding the variable manually.

• Sectioning commands are now shown in a larger font on display devices which support such fontification. The variable font-latex-title-fontify can be customized to restore the old appearance, i.e. the usage of a different color instead of a change in size.

• Support for alphanum.sty, beamer.cls, booktabs.sty, captcont.sty, emp.sty,
paralist.sty, subfigure.sty and units.sty/nicefrac.sty was added. Credits

go to the authors mentioned in the respective AUCTEX style files.

• Inserting graphics with C-c RET \includegraphics RET was improved. See the vari-

able LaTeX-includegraphics-options-alist.

• If LaTeX-default-position is nil, don’t prompt for position arguments in Tabularlike environments, see Section 4.4.4 [Tabular-like], page 26.

• Completion for available packages when using C-c RET \usepackage RET was improved on systems using the kpathsea library.

• The commenting functionality was fixed. The separate functions for commenting and uncommenting were unified in one function for paragraphs and regions respectively which do both.

• Syntax highlighting can be customized to fontify quotes delimited by either >>German<< or <<French>> quotation marks by changing the variable font-latex-quotes.

• Certain TEX/LaTEX keywords for functions, references, variables and warnings will now be fontified specially. You may add your own keywords by customizing the variables

font-latex-match-function-keywords, font-latex-match-reference-keywords,
font-latex-match-variable-keywords and font-latex-match-warning-keywords.

• If you include the style files ‘german’ or ‘ngerman’ in a document (directly or via the ‘babel’ package), you should now customize LaTeX-german-open-quote, LaTeX-

german-close-quote and LaTeX-german-quote-after-quote instead of TeX-open-
quote, TeX-close-quote and TeX-quote-after-quote if you want to influence the

type of quote insertion.

• Upon viewing an output file, the right viewer and command line options for it are now determined automatically by looking at the extension of the output file and certain options used in the source file. The behavior can be adapted or extended respectively by customizing the variable TeX-output-view-style.

• You can control whether TeX-insert-macro (C-c RET) ask for all optional arguments by customizing the variable TeX-insert-macro-default-style, Section 5.2 [Completion], page 29.

• TeX-run-discard is now able to completely detach a process that it started.

• The build process was enhanced and is now based on autoconf making installing AUCTEX a mostly automatic process. See Chapter 2 [Installation], page 5 and Section 2.7 [Installation under MS Windows], page 11 for details.

News in 11.14

• Many more LaTeX and LaTeX2e commands are supported. Done by Masayuki Ataka <ataka@milk.freemail.ne.jp>

News in 11.12

• Support for the KOMA-Script classes. Contributed by Mark Trettin <Mark.Trettin@gmx.de>.

News in 11.11

• Support for ‘prosper.sty’, see http://prosper.sourceforge.net/. Contributed by

Phillip Lord <p.lord@russet.org.uk>.

News in 11.10

• comment-region now inserts %% by default. Suggested by "Davide G. M. Salvetti" <salve@debian.org>.

News in 11.06

• You can now switch between using the ‘font-latex’ (all emacsen), the ‘tex-font’ (Emacs 21 only) or no special package for font locking. Customize TeX-install-

font-lock for this.

News in 11.04

• Now use -t landscape by default when landscape option appears. Suggested by Erik Frisk <frisk@isy.liu.se>.

News in 11.03

• Use ‘tex-fptex.el’ for fpTeX support. Contributed by Fabrice Popineau <Fabrice.Popineau@supelec.fr>.

News in 11.02

• New user option LaTeX-top-caption-list specifies environments where the caption should go at top. Contributed by ataka@milk.freemail.ne.jp (Masayuki Ataka).

• Allow explicit dimensions in ‘graphicx.sty’. Contributed by ataka@milk.freemail.ne.jp (Masayuki Ataka).

• Limited support for ‘verbatim.sty’. Contributed by ataka@milk.freemail.ne.jp (Masayuki Ataka).

• Better support for asmmath items. Patch by ataka@milk.freemail.ne.jp (Masayuki Ataka).

• More accurate error parsing. Added by David Kastrup <David.Kastrup@t-online.de>.

News in 11.01

• Bug fixes.

Older versions

See the file ‘history.texi’ for older changes.

Appendix C Future Development

The following sections describe future development of AUCTEX. Besides mid-term goals, bug reports and requests we cannot fix or honor right away are being gathered here. If you have some time for Emacs Lisp hacking, you are encouraged to try to provide a solution to one of the following problems. If you don’t know Lisp, you may help us to improve the documentation. It might be a good idea to discuss proposed changes on the mailing list of AUCTEX first.

C.1 Mid-term Goals

• Integration of preview-latex into AUCTEX As of AUCTEX 11.81 preview-latex is a part of AUCTEX in the sense that the installation routines were merged and preview-latex is being packaged with AUCTEX. Further integration will happen at the backend. This involves folding of error parsing and task management of both packages which will ease development efforts and avoid redundant work.

• More flexible option and command handling The current state of command handling with TeX-command-list is not very flexible because there is no distinction between executables and command line options to be passed to them. Customization of TeX-command-list by the user will interfere with updates of AUCTEX.

• Error help catalogs Currently, the help for errors is more or less hardwired into ‘tex.el’. For supporting error help in other languages, it would be sensible to instead arrange error messages in language-specific files, make a common info file from all such catalogs in a given language and look the error texts up in an appropriate index. The user would then specify a preference list of languages, and the errors would be looked up in the catalogs in sequence until they were identified.

• Combining ‘docTeX’ with RefTeX Macro cross references should also be usable for document navigation using RefTeX.

C.2 Wishlist

• Documentation lookup for macros A parser could gather information about which macros are defined in which LaTEX packages and store the information in a hashtable which can be used in a backend for TeX-doc in order to open the matching documentation for a given macro. The information could also be used to insert an appropriate ‘\usepackage’ statement if the user tries to insert a macro for which the respective package has not been requested yet.

• Spell checking of macros A special ispell dictionary for macros could be nice to have.

• Quick error overviews An error overview window (extract from the log file with just the error lines, clickable like a “grep” buffer) and/or fringe indicators for errors in the main text would be nice.

• A math entry grid A separate frame with a table of math character graphics to click on in order to insert the respective sequence into the buffer (cf. the “grid” of x-symbol).

• Crossreferencing support It would be nice if you could index process your favorite collection of ‘.dtx’ files (such as the LaTeX source), just call a command on arbitrary control sequence, and get either the DVI viewer opened right at the definition of that macro (using Source Specials), or the source code of the ‘.dtx’ file.

• Better plain TeX support For starters, LaTeX-math-mode is not very LaTEX-specific in the first place, and similar holds for indentation and formatting.

• Poor man’s Source Specials In particular in PDF mode (and where Source Specials cause problems), alternatives would be desirable. One could implement inverse search by something like Heiko Oberdiek’s ‘vpe.sty’, and forward search by using the ‘.aux’ file info to correlate labels in the text (possibly in cooperation with RefTEX) with previewer pages. In AUCTEX 11.83, support for forward search with PDF files was added. Currently this only works if you use the pdfsync LaTEX package and xpdf as your PDF viewer. See Section 7.2 [Viewing], page 46.

• Page count when compiling should (optionally) go to modeline of the window where the compilation command was invoked, instead of the output window. Suggested by Karsten Tinnefeld <tinnefeld@irb.informatik.uni-dortmund.de>.

• Command to insert a macrodefinition in the preamble, without moving point from the current location. Suggested by "Jeffrey C. Ely" <ely@nwu.edu>.

• A database of all commands defined in all stylefiles. When a command or environment gets entered that is provided in one of the styles, insert the appropriate \usepackage in the preamble.

• A way to add and overwrite math mode entries in style files, and to decide where they should be. Suggested by Remo Badii <Remo.Badii@psi.ch>.

• Create template for (first) line of tabular environment.

• I think prompting for the master is the intended behaviour. It corresponds to a ‘shared’ value for TeX-master. There should probably be a ‘none’ value which wouldn’t query for the master, but instead disable all features that relies on TeX-master. This default value for TeX-master could then be controled with mapping based on the extension.

• Multiple argument completion for ‘\bibliography’. In general, I ought to make , special for these kind of completions.

• Suggest ‘makindex’ when appropriate.

• Use index files (when available) to speed up C-c C-m include ^{⟨}\underline{\overline{RET}}^{⟩}.

• Option not to calculate very slow completions like for C-c C-m include ^{⟨}\underline{\overline{RET}}^{⟩}.

• Font menu should be created from TeX-font-list.

• Installation procedure written purely in emacs lisp.

• Included PostScript files should also be counted as part of the document.

• The parser should catch warnings about undefined crossreferences. Suggested by

Richard Hirsch ‘i3080501@ws.rz.tu-bs.de’.

• A nice hierarchical by-topic organization of all officially documented LaTeX macros, available from the menu bar.

• TeX-command-default should be set from the master file, if not set locally. Suggested

by Peter Whaite ‘<peta@cim.mcgill.ca>’.

• Make AUCTEX work with ‘crypt++’. Suggested by Chris Moore

‘<Chris.Moore@src.bae.co.uk>’.

• Make AUCTEX work with ‘longlines’. This would also apply to preview-latex, though it might make sense to unify error processing before attempting this.

• The ‘Spell’ command should apply to all files in a document. Maybe it could try to restrict to files that have been modified since last spell check? Suggested by Ravinder

Bhumbla ‘<rbhumbla@ucsd.edu>’.

• Make ^{⟨}\underline{\overline{.}}^{⟩} check for abbreviations and sentences ending with capital letters.

• Use Emacs 19 minibuffer history to choose between previewers, and other stuff. Sug-

gested by John Interrante ‘<interran@uluru.Stanford.EDU>’.

• Make features.

A new command TeX-update (C-c C-u) could be used to create an up-to-date dvi file by repeatedly running BibTEX, MakeIndex and (La)TEX, until an error occurs or we are done. An alternative is to have an ‘Update’ command that ensures the ‘dvi’ file is up to date. This could be called before printing and previewing.

• Documentation of variables that can be set in a style hook.

We need a list of what can safely be done in an ordinary style hook. You can not set a variable that AUCTEX depends on, unless AUCTEX knows that it has to run the style hooks first. Here is the start of such a list.

LaTeX-add-environments
TeX-add-symbols
LaTeX-add-labels
LaTeX-add-bibliographies
LaTeX-largest-level

• Completion for counters and sboxes.

• Outline should be (better) supported in TEX mode.

At least, support headers, trailers, as well as TeX-outline-extra.

• TeX-header-start and TeX-trailer-end.

We might want these, just for fun (and outlines)

• Plain TEX and LaTEX specific header and trailer expressions.

We should have a way to globally specify the default value of the header and trailer regexps.

• Get closer to original TeX-mode keybindings.

A third initialization file (‘tex-mode.el’) containing an emulator of the standard TeXmode would help convince some people to change to AUCTEX.

• Make TeX-next-error parse ahead and store the results in a list, using markers to remember buffer positions in order to be more robust with regard to line numbers and changed files. This is what next-error does. (Or did, until Emacs 19).

• Finish the Texinfo mode. For one thing, many Texinfo mode commands do not accept braces around their arguments.

• Hook up the letter environment with ‘bbdb.el’.

C.3 Bugs

• The parsed files and style hooks for ‘example.dtx’, ‘example.sty’, ‘example.drv’ and ‘example.bib’ all clash. Bad.

• C-c ` should always stay in the current window, also when it finds a new file.

• Do not overwrite emacs warnings about existing auto-save files when loading a new file.

• Maybe the regexp for matching a TeX symbol during parsing should be

  ‘"\\\\\\([a-zA-Z]+\\|.\\)"’ — ‘<thiemann@informatik.uni-tuebingen.de>’ Pe-

ter Thiemann.

• AUCTEX should not parse verbatim environments.

• Make ‘`’ check for math context in LaTeX-math-mode. and simply self insert if not in a math context.

• Make TeX-insert-dollar more robust. Currently it can be fooled by ‘\mbox’’es and escaped double dollar for example.

• Correct indentation for tabular, tabbing, table, math, and array environments.

• Syntax highlighting of LaTEX constructs spanning more than one line sometimes stops in the middle of the construct. Highlighting can get lost during typing. (XEmacs only)

• No syntactic font locking of verbatim macros and environments. (XEmacs only)

• Font locking inside of verbatim macros and environments is not inhibited. This may result in syntax highlighting of unbalanced dollar signs and the like spilling out of the verbatim content. (XEmacs only)

• Folding of LaTEX constructs spanning more than one line may result in overfull lines. (XEmacs only)

Appendix D Frequently Asked Questions

1. Something is not working correctly. What should I do? Well, you might have guessed it, the first place to look is in the available documentation packaged with AUCTEX. This could be the release notes (in the ‘RELEASE’ file) or the news section of the manual in case you are experiencing problems after an upgrade, the ‘INSTALL’ file in case you are having problems with the installation, the section about bugs in the manual in case you encountered a bug or the relevant sections in the manual for other related problems. If this did not help, you can send a bug report to the AUCTEX bug reporting list by using the command M-x TeX-submit-bug-report RET. But before you do this, you can try to get more information about the problem at hand which might also help you locate the cause of the error yourself. First, you can try to generate a so-called backtrace which shows functions involved in a program error. In order to do this, start Emacs with the command line ‘emacs --debug-init’ and/or put the line

     (setq debug-on-error t)

as the first line into your init file. XEmacs users might want to add (setq stacktrace-on-error t) as well. After Emacs has started, you can load a file which triggers the error and a new window should pop up showing the backtrace. If you get such a backtrace, please include it in the bug report. Second, you can try to figure out if something in your personal or site configuration triggers the error by starting Emacs without such customizations. You can do this by invoking Emacs with the command line ‘emacs -q -no-site-file’. Once Emacs is running, copy the line

     (load "auctex.el" nil t t)

into the ‘*scratch*’ buffer and type M-x eval-buffer RET. This makes sure that AUCTEX will be used for the file types it supports. After you have done so, you can load the file triggering the error. If everything is working now, you know that you have to search either in the site configuration file or your personal init file for statements related to the problem.

2. What versions of Emacs and XEmacs are supported? AUCTEX was tested with Emacs 21 and XEmacs 21.4.15. Older versions may work but are unsupported. Older versions of XEmacs might possibly made to work by updating the ‘xemacs-base’ package through the XEmacs package system. If you are looking for a recommendation, it would appear that the smoothest working platform on all operating systems at the current point of time would be Emacs 22.1. At the time of this writing, however, it has not been released and is still under development. The quality of the development version is quite solid, so we recommend giving it a try. With a developer version, of course, you have to be prepared to update in case you managed to get your snapshot at a bad time. The second best choice would be the latest released Emacs 21.4. However, Unicode support is less good, there is no version for the popular GTK toolkit, and the native versions for Windows and MacOS don’t offer toolbar and preview-latex support.

Our success with XEmacs has been less than convincing. Under the Windows operating system, nominally the only option for a released, stable Emacs variant supporting toolbars and preview-latex would be XEmacs 21.4. However, code for core functionality like formatting and syntax highlighting tends to be different and often older than even Emacs 21.4, and Unicode support as delivered is problematic at best, missing on Windows. Both AUCTEX and XEmacs developers don’t hear much from active users of the combination. Partly for that reason, problems tend to go unnoticed for long amounts of time and are often found, if at all, after releases. No experiences or recommendations can be given for beta or developer versions of XEmacs.

3. Why doesn’t the completion, style file, or multi-file stuff work?

It must be enabled first, insert this in your init file:

     (setq-default TeX-master nil)
     (setq TeX-parse-self t)
     (setq TeX-auto-save t)

Read also the chapters about parsing and multifile documents in the manual.

4. Why doesn’t TeX-save-document work?

TeX-check-path has to contain "./" somewhere.

5. Why is the information in ‘foo.tex’ forgotten when I save ‘foo.bib’? For various reasons, AUCTEX ignores the extension when it stores information about a file, so you should use unique base names for your files. E.g. rename ‘foo.bib’ to

‘foob.bib’.

6. Why doesn’t AUCTEX signal when processing a document is done? If the message in the minibuffer stays "Type ‘C-c C-l’ to display results of compilation.", you probably have a misconfiguration in your init file (‘.emacs’, ‘init.el’ or similar). To track this down either search in the ‘*Messages*’ buffer for an error message or put (setq debug-on-error t) as the first line into your init file, restart Emacs and open a LaTEX file. Emacs will complain loudly by opening a debugging buffer as soon as an error occurs. The information in the debugging buffer can help you find the cause of the error in your init file.

7. What does AUC stand for?

AUCTEX came into being at Aalborg University in Denmark. Back then the Danish name of the university was Aalborg Universitetscenter; AUC for short.

Key Index 90

Key Index

"

" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

$

$ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

C

C-c % . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 C-c ; . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 C-c ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 C-c ] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 C-c ^. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 C-c _. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 C-c ` . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 C-c { . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 C-c ~ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 C-c C-b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 C-c C-c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 C-c C-d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 C-c C-e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 C-c C-f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 C-c C-f C-b. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 22 C-c C-f C-c. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 22 C-c C-f C-e. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 22 C-c C-f C-f. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 22 C-c C-f C-i. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 22 C-c C-f C-r. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 22 C-c C-f C-s. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 22 C-c C-f C-t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 22 C-c C-k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 C-c C-l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 C-c C-m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 C-c C-n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 C-c C-o b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 C-c C-o C-b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 C-c C-o C-c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 C-c C-o C-e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 C-c C-o C-f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 C-c C-o C-m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 C-c C-o C-o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 C-c C-o C-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 C-c C-o C-r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 C-c C-o i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 C-c C-o p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 C-c C-o r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 C-c C-q C-e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 C-c C-q C-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 C-c C-q C-r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 C-c C-q C-s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 C-c C-r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 C-c C-s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 C-c C-t C-b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 C-c C-t C-i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 C-c C-t C-o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 C-c C-t C-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 C-c C-t C-r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 C-c C-t C-s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 C-c C-t C-w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 C-c C-v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 C-c ⟨\underline{\overline{LFD}}⟩. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 C-j . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

L

⟨\underline{\overline{LFD}}⟩. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

M

M-q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 M-⟨\underline{\overline{TAB}}⟩. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

T

⟨\underline{\overline{TAB}}⟩. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

Function Index 91

Function Index

L

LaTeX-add-bibliographies . . . . . . . . . . . . . . . . . . 66
LaTeX-add-environments . . . . . . . . . . . . . . . . . . . . 66

LaTeX-add-labels . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

LaTeX-close-environment . . . . . . . . . . . . . . . . . . . 25

LaTeX-env-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 LaTeX-environment . . . . . . . . . . . . . . . . . . . . . . . . . . 25

LaTeX-fill-environment . . . . . . . . . . . . . . . . . . . . 34

LaTeX-fill-paragraph. . . . . . . . . . . . . . . . . . . . . . . 34 LaTeX-fill-region . . . . . . . . . . . . . . . . . . . . . . . . . . 34 LaTeX-fill-section . . . . . . . . . . . . . . . . . . . . . . . . . 34 LaTeX-indent-line . . . . . . . . . . . . . . . . . . . . . . . . . . 31

LaTeX-insert-environment . . . . . . . . . . . . . . . . . . 66

LaTeX-insert-item . . . . . . . . . . . . . . . . . . . . . . . . . . 26 LaTeX-math-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 LaTeX-section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 LaTeX-section-heading . . . . . . . . . . . . . . . . . . . . . 24 LaTeX-section-label. . . . . . . . . . . . . . . . . . . . . . . . 24 LaTeX-section-section . . . . . . . . . . . . . . . . . . . . . 24 LaTeX-section-title. . . . . . . . . . . . . . . . . . . . . . . . 24 LaTeX-section-toc . . . . . . . . . . . . . . . . . . . . . . . . . . 24

T

TeX-add-style-hook . . . . . . . . . . . . . . . . . . . . . . . . . 62 TeX-add-symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 TeX-arg-cite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 TeX-arg-conditional. . . . . . . . . . . . . . . . . . . . . . . . 63 TeX-arg-coordinate . . . . . . . . . . . . . . . . . . . . . . . . . 65 TeX-arg-corner . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 TeX-arg-counter . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 TeX-arg-define-cite. . . . . . . . . . . . . . . . . . . . . . . . 64

TeX-arg-define-counter . . . . . . . . . . . . . . . . . . . . 64
TeX-arg-define-environment . . . . . . . . . . . . . . . . 64

TeX-arg-define-label. . . . . . . . . . . . . . . . . . . . . . . 64 TeX-arg-define-macro. . . . . . . . . . . . . . . . . . . . . . . 64

TeX-arg-define-savebox . . . . . . . . . . . . . . . . . . . . 64

TeX-arg-environment. . . . . . . . . . . . . . . . . . . . . . . . 64 TeX-arg-eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 TeX-arg-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 TeX-arg-free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 TeX-arg-input-file . . . . . . . . . . . . . . . . . . . . . . . . . 64 TeX-arg-label . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 TeX-arg-literal . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 TeX-arg-lr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 TeX-arg-macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 TeX-arg-pagestyle . . . . . . . . . . . . . . . . . . . . . . . . . . 64 TeX-arg-pair . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 TeX-arg-savebox . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 TeX-arg-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 TeX-arg-tb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 TeX-arg-verb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 TeX-auto-generate . . . . . . . . . . . . . . . . . . . . . . . . . . 60 TeX-clean . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 TeX-command-buffer . . . . . . . . . . . . . . . . . . . . . . . . . 45 TeX-command-master . . . . . . . . . . . . . . . . . . . . . . . . . 44 TeX-command-region . . . . . . . . . . . . . . . . . . . . . . . . . 44

TeX-comment-or-uncomment-paragraph . . . . . . . 30
TeX-comment-or-uncomment-region . . . . . . . . . . . 30

TeX-complete-symbol. . . . . . . . . . . . . . . . . . . . . . . . 29 TeX-doc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 TeX-electric-macro . . . . . . . . . . . . . . . . . . . . . . . . . 30 TeX-fold-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

TeX-fold-clearout-buffer . . . . . . . . . . . . . . . . . . 41
TeX-fold-clearout-item . . . . . . . . . . . . . . . . . . . . 42
TeX-fold-clearout-paragraph . . . . . . . . . . . . . . . 42
TeX-fold-clearout-region . . . . . . . . . . . . . . . . . . 42

TeX-fold-comment . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 TeX-fold-dwim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 TeX-fold-env . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 TeX-fold-macro . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 TeX-fold-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 TeX-fold-paragraph . . . . . . . . . . . . . . . . . . . . . . . . . 41 TeX-fold-region . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 TeX-font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 TeX-header-end . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 TeX-home-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 TeX-insert-braces . . . . . . . . . . . . . . . . . . . . . . . . . . 22 TeX-insert-dollar . . . . . . . . . . . . . . . . . . . . . . . . . . 22 TeX-insert-macro . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 TeX-insert-quote . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 TeX-interactive-mode. . . . . . . . . . . . . . . . . . . . . . . 46 TeX-kill-job . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 TeX-master-file-ask. . . . . . . . . . . . . . . . . . . . . . . . 52 TeX-next-error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 TeX-normal-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 TeX-Omega-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 TeX-PDF-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 TeX-pin-region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

TeX-recenter-output-buffer . . . . . . . . . . . . . . . . 49

TeX-save-document . . . . . . . . . . . . . . . . . . . . . . . . . . 52

TeX-source-specials-mode . . . . . . . . . . . . . . . . . . 46
TeX-toggle-debug-bad-boxes . . . . . . . . . . . . . . . . 48
TeX-toggle-debug-warnings . . . . . . . . . . . . . . . . . 48

TeX-view. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

Variable Index 92

Variable Index

C

ConTeXt-clean-intermediate-suffixes . . . . . . 49
ConTeXt-clean-output-suffixes . . . . . . . . . . . . . 49

D

docTeX-clean-intermediate-suffixes . . . . . . . 49
docTeX-clean-output-suffixes . . . . . . . . . . . . . . 49

F

font-latex-deactivated-keyword-classes . . . 38
font-latex-do-multi-line . . . . . . . . . . . . . . . . . . 40
font-latex-fontify-script . . . . . . . . . . . . . . . . . 39
font-latex-fontify-sectioning . . . . . . . . . . . . . 37
font-latex-match-bold-command-keywords . . . 37
font-latex-match-bold-declaration-keywords

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

font-latex-match-function-keywords . . . . . . . 36
font-latex-match-italic-command-keywords

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

font-latex-match-italic-declaration-

keywords . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

font-latex-match-math-command-keywords . . . 37
font-latex-match-reference-keywords . . . . . . 36
font-latex-match-sectioning-0-keywords . . . 37
font-latex-match-sectioning-1-keywords . . . 37
font-latex-match-sectioning-2-keywords . . . 37
font-latex-match-sectioning-3-keywords . . . 37
font-latex-match-sectioning-4-keywords . . . 37
font-latex-match-sectioning-5-keywords . . . 37
font-latex-match-slide-title-keywords . . . . 37
font-latex-match-textual-keywords. . . . . . . . . 36
font-latex-match-type-command-keywords . . . 37
font-latex-match-type-declaration-keywords

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

font-latex-match-variable-keywords . . . . . . . 36
font-latex-match-warning-keywords. . . . . . . . . 36

font-latex-quotes . . . . . . . . . . . . . . . . . . . . . . . . . . 39

font-latex-script-display . . . . . . . . . . . . . . . . . 39
font-latex-sectioning-0-face . . . . . . . . . . . . . . 37
font-latex-sectioning-1-face . . . . . . . . . . . . . . 37
font-latex-sectioning-2-face . . . . . . . . . . . . . . 37
font-latex-sectioning-3-face . . . . . . . . . . . . . . 37
font-latex-sectioning-4-face . . . . . . . . . . . . . . 37
font-latex-sectioning-5-face . . . . . . . . . . . . . . 37
font-latex-slide-title-face . . . . . . . . . . . . . . . 37
font-latex-user-keyword-classes . . . . . . . . . . . 38

J

japanese-LaTeX-command-default . . . . . . . . 57, 58
japanese-LaTeX-default-style. . . . . . . . . . . 57, 58
japanese-TeX-command-default . . . . . . . . . . . . . . 57

L

LaTeX-amsmath-label. . . . . . . . . . . . . . . . . . . . . . . . 26

LaTeX-auto-label-regexp-list . . . . . . . . . . . . . . 54
LaTeX-auto-minimal-regexp-list . . . . . . . . . . . . 54
LaTeX-auto-regexp-list . . . . . . . . . . . . . . . . . . . . 54

LaTeX-babel-hyphen . . . . . . . . . . . . . . . . . . . . . . . . . 57

LaTeX-babel-hyphen-after-hyphen . . . . . . . . . . . 57
LaTeX-babel-hyphen-language-alist. . . . . . . . . 57
LaTeX-clean-intermediate-suffixes. . . . . . . . . 49
LaTeX-clean-output-suffixes . . . . . . . . . . . . . . . 49
LaTeX-csquotes-close-quote . . . . . . . . . . . . . . . . 21
LaTeX-csquotes-open-quote . . . . . . . . . . . . . . . . . 21
LaTeX-csquotes-quote-after-quote. . . . . . . . . . 21
LaTeX-default-environment . . . . . . . . . . . . . . . . . 25

LaTeX-default-format. . . . . . . . . . . . . . . . . . . . . . . 27

LaTeX-default-position . . . . . . . . . . . . . . . . . . . . 27

LaTeX-enable-toolbar. . . . . . . . . . . . . . . . . . . . . . . 44 LaTeX-eqnarray-label. . . . . . . . . . . . . . . . . . . . . . . 25 LaTeX-equation-label. . . . . . . . . . . . . . . . . . . . . . . 25 LaTeX-figure-label . . . . . . . . . . . . . . . . . . . . . . . . . 26

LaTeX-fill-break-at-separators . . . . . . . . . . . . 34
LaTeX-fill-break-before-code-comments . . . . 34

LaTeX-float . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

LaTeX-indent-environment-check . . . . . . . . . . . . 31
LaTeX-indent-environment-list . . . . . . . . . 31, 32

LaTeX-indent-level . . . . . . . . . . . . . . . . . . . . . 31, 32 LaTeX-item-indent . . . . . . . . . . . . . . . . . . . . . . 31, 32

LaTeX-math-abbrev-prefix . . . . . . . . . . . . . . . . . . 28

LaTeX-math-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

LaTeX-math-menu-unicode . . . . . . . . . . . . . . . . . . . 28
LaTeX-paragraph-commands . . . . . . . . . . . . . . . . . . 33

LaTeX-section-hook . . . . . . . . . . . . . . . . . . . . . 23, 24 LaTeX-section-label . . . . . . . . . . . . . . . . . . . . 23, 24

LaTeX-syntactic-comments . . . . . . . . . . . . . . . 31, 32

LaTeX-table-label . . . . . . . . . . . . . . . . . . . . . . . . . . 26

LaTeX-top-caption-list . . . . . . . . . . . . . . . . . . . . 26
LaTeX-verbatim-environments . . . . . . . . . . . . . . . 39
LaTeX-verbatim-macros-with-braces. . . . . . . . . 39
LaTeX-verbatim-macros-with-delims. . . . . . . . . 39

P

plain-TeX-auto-regexp-list . . . . . . . . . . . . . . . . 54
plain-TeX-clean-intermediate-suffixes . . . . 49
plain-TeX-clean-output-suffixes . . . . . . . . . . . 49
plain-TeX-enable-toolbar . . . . . . . . . . . . . . . . . . 44

T

TeX-auto-cleanup-hook . . . . . . . . . . . . . . . . . . . . . 68

TeX-auto-empty-regexp-list . . . . . . . . . . . . . . . . 54
TeX-auto-full-regexp-list . . . . . . . . . . . . . . . . . 54

TeX-auto-global . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 TeX-auto-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

Variable Index 93

TeX-auto-parse-length . . . . . . . . . . . . . . . . . . . . . 54 TeX-auto-prepare-hook . . . . . . . . . . . . . . . . . . . . . 68 TeX-auto-private . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 TeX-auto-regexp-list . . . . . . . . . . . . . . . . . . . 54, 68 TeX-auto-save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 TeX-auto-untabify . . . . . . . . . . . . . . . . . . . . . . . . . . 53

TeX-brace-indent-level . . . . . . . . . . . . . . . . . . . . 32

TeX-check-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 TeX-clean-confirm . . . . . . . . . . . . . . . . . . . . . . . . . . 50 TeX-close-quote . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 TeX-command-default. . . . . . . . . . . . . . . . . . . . . . . . 45 TeX-command-list . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 TeX-default-macro . . . . . . . . . . . . . . . . . . . . . . . . . . 29 TeX-default-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 TeX-display-help . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 TeX-DVI-via-PDFTeX . . . . . . . . . . . . . . . . . . . . . . . . . 46 TeX-electric-escape. . . . . . . . . . . . . . . . . . . . . . . . 30

TeX-electric-sub-and-superscript. . . . . . . . . . 29

TeX-file-recurse . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

TeX-fold-env-spec-list . . . . . . . . . . . . . . . . . . . . 42
TeX-fold-force-fontify . . . . . . . . . . . . . . . . . . . . 41
TeX-fold-help-echo-max-length . . . . . . . . . . . . . 43
TeX-fold-macro-spec-list . . . . . . . . . . . . . . . . . . 42
TeX-fold-preserve-comments . . . . . . . . . . . . . . . . 41

TeX-fold-type-list . . . . . . . . . . . . . . . . . . . . . . . . . 41

TeX-fold-unspec-env-display-string . . . . . . . 42
TeX-fold-unspec-macro-display-string . . . . . 42
TeX-fold-unspec-use-name . . . . . . . . . . . . . . . . . . 42

TeX-font-list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 TeX-header-end. . . . . . . . . . . . . . . . . . . . . . . . . . 44, 45 TeX-ignore-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 TeX-insert-braces . . . . . . . . . . . . . . . . . . . . . . . . . . 30

TeX-insert-macro-default-style . . . . . . . . . . . . 29

TeX-install-font-lock . . . . . . . . . . . . . . . . . . . . . 35 TeX-interactive-mode. . . . . . . . . . . . . . . . . . . . . . . 46 TeX-language-cz-hook. . . . . . . . . . . . . . . . . . . . . . . 56 TeX-language-de-hook. . . . . . . . . . . . . . . . . . . . . . . 56 TeX-language-dk-hook. . . . . . . . . . . . . . . . . . . . . . . 56 TeX-language-it-hook. . . . . . . . . . . . . . . . . . . . . . . 56 TeX-language-nl-hook. . . . . . . . . . . . . . . . . . . . . . . 56 TeX-language-pl-hook. . . . . . . . . . . . . . . . . . . . . . . 56 TeX-language-sk-hook. . . . . . . . . . . . . . . . . . . . . . . 56 TeX-language-sv-hook. . . . . . . . . . . . . . . . . . . . . . . 56 TeX-macro-global . . . . . . . . . . . . . . . . . . . . . . . 16, 59 TeX-macro-private . . . . . . . . . . . . . . . . . . . . . . . . . . 60 TeX-master . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44, 51 TeX-newline-function . . . . . . . . . . . . . . . . . . . 31, 32 TeX-Omega-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 TeX-one-master . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 TeX-open-quote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 TeX-outline-extra . . . . . . . . . . . . . . . . . . . . . . . . . . 43 TeX-output-view-style . . . . . . . . . . . . . . . . . . . . . 47 TeX-parse-self . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 TeX-PDF-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 TeX-quote-after-quote . . . . . . . . . . . . . . . . . . . . . 21

                                                TeX-quote-language-alist . . . . . . . . . . . . . . . . . . 56

TeX-region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44, 45 TeX-save-query . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

                                                TeX-source-specials-mode . . . . . . . . . . . . . . . . . . 46
                                                TeX-source-specials-view-start-server . . . . 47

TeX-style-global . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 TeX-style-local . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 TeX-style-path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 TeX-style-private . . . . . . . . . . . . . . . . . . . . . . . . . . 60 TeX-trailer-start . . . . . . . . . . . . . . . . . . . . . . 44, 45 TeX-view-style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

                                                Texinfo-clean-intermediate-suffixes . . . . . . 49
                                                Texinfo-clean-output-suffixes . . . . . . . . . . . . . 49

Concept Index 94

Concept Index

.

‘.emacs’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

\

‘\begin’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 \chapter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 23 \cite, completion of . . . . . . . . . . . . . . . . . . . . . . . . . . 30 \emph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 22 ‘\end’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 \include. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 \label . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 23 \label, completion . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 \ref, completion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 \section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 23 \subsection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 23 \textbf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 22 \textit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 22 \textrm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 22 \textsc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 22 \textsf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 22 \textsl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 22 \texttt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 22

A

Abbreviations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Adding a style hook. . . . . . . . . . . . . . . . . . . . . . . . . . 62 Adding bibliographies . . . . . . . . . . . . . . . . . . . . . . . . 66 Adding environments . . . . . . . . . . . . . . . . . . . . . . . . 65 Adding labels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 Adding macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Adding other information . . . . . . . . . . . . . . . . . . . . 66 Adding to PATH in Windows . . . . . . . . . . . . . . . . . . 12 Advanced features . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 amsmath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 ANSI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Arguments to TeX macros . . . . . . . . . . . . . . . . . . . . 29 ASCII pTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55, 57 ‘auto’ directories. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 Auto-Reveal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 Automatic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 Automatic Customization . . . . . . . . . . . . . . . . . . . . 59 Automatic Parsing . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Automatic updating style hooks . . . . . . . . . . . . . . 60

B

Bad boxes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 Bibliographies, adding. . . . . . . . . . . . . . . . . . . . . . . . 66 Bibliography . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 bibliography, completion . . . . . . . . . . . . . . . . . . . . . 30 BibTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 BibTeX, completion. . . . . . . . . . . . . . . . . . . . . . . . . . 30 ‘book.el’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Braces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Brackets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

C

Changing font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Changing the parser . . . . . . . . . . . . . . . . . . . . . . . . . 66 Chapters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 23 Character set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Checking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 ChinaTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 chktex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 citations, completion of . . . . . . . . . . . . . . . . . . . . . . 30 cite, completion of . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 CJK language . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 CJK-LaTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Cleaning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Completion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Controlling the output . . . . . . . . . . . . . . . . . . . . . . . 49 Copying. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 CTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Current file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Customization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Customization, personal . . . . . . . . . . . . . . . . . . . . . . 16 Customization, site . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Czech . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

D

Danish . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 Default command . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Defining bibliographies in style hooks. . . . . . . . . . 66 Defining environments in style hooks . . . . . . . . . . 65 Defining labels in style hooks . . . . . . . . . . . . . . . . . 66 Defining macros in style hooks . . . . . . . . . . . . . . . . 62 Defining other information in style hooks . . . . . . 66 Deleting fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19, 22 Denmark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Descriptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Display math mode . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 Documents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Documents with multiple files. . . . . . . . . . . . . . . . . 51 Dollars. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Double quotes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Dutch. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

E

Enumerates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Environments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Environments, adding . . . . . . . . . . . . . . . . . . . . . . . . 65 Eqnarray. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Equation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Equations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 Europe. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 European Characters. . . . . . . . . . . . . . . . . . . . . . . . . 55 Example of a style file. . . . . . . . . . . . . . . . . . . . . . . . 62 Expansion. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 External Commands . . . . . . . . . . . . . . . . . . . . . . . . . 44 Extracting TeX symbols . . . . . . . . . . . . . . . . . . . . . . 59

F

Faces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 FDL, GNU Free Documentation License . . . . . . . 69 Figure environment . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Figures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 File Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 Filling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Finding errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Finding the current file. . . . . . . . . . . . . . . . . . . . . . . 49 Finding the master file . . . . . . . . . . . . . . . . . . . . . . . 49 Floats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Folding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40, 43 Font Locking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Font macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 font-latex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Formatting . . . . . . . . . . . . . . . . . . . . . . . . . . . 31, 32, 44 Forward search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 Free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Free software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

G

General Public License . . . . . . . . . . . . . . . . . . . . . . . . 2 Generating symbols . . . . . . . . . . . . . . . . . . . . . . . . . . 59 Germany . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Global directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 Global macro directory . . . . . . . . . . . . . . . . . . . . . . . 59 Global style hook directory . . . . . . . . . . . . . . . . . . . 59 Global TeX macro directory . . . . . . . . . . . . . . . . . . 59 GPL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

H

Header. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 Hide Macros. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 HLaTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Holland . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

I

Including. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Indentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 Indenting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 Indexing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Initialization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Inputing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Internationalization . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Inverse search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 ISO 8859 Latin 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 ISO 8859 Latin 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 ‘iso-cvt.el’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 ispell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Italy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Itemize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

J

Japan. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Japanese . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 jLaTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 jTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55, 57

K

Killing a process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 kTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

L

Label prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25, 26 Labels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25, 26 Labels, adding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 labels, completion of . . . . . . . . . . . . . . . . . . . . . . . . . 30 lacheck. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Language Support . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 LaTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Latin 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Latin 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Literature . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Local style directory . . . . . . . . . . . . . . . . . . . . . . . . . 60 Local style hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 Local Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

M

Macro arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Macro completion. . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Macro expansion. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 ‘macro.el’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 ‘macro.tex’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 Macros, adding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 makeindex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Making a bibliography . . . . . . . . . . . . . . . . . . . . . . . 44 Making an index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Many Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Master file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 51 Matching dollar signs . . . . . . . . . . . . . . . . . . . . . . . . 21 Math mode delimiters . . . . . . . . . . . . . . . . . . . . . . . . 21 Mathematics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 MULE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55, 57 MULE-UCS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Multifile Documents . . . . . . . . . . . . . . . . . . . . . . . . . 51 Multiple Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

N

National letters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Next error. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 Nippon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 NTT jTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55, 57

O

Omega . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 Other information, adding . . . . . . . . . . . . . . . . . . . . 66 Outlining . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40, 43 Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Overfull boxes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 Overview. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

P

Parsing errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 Parsing LaTeX errors . . . . . . . . . . . . . . . . . . . . . . . . 48 Parsing new macros . . . . . . . . . . . . . . . . . . . . . . . . . . 66 Parsing TeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53, 59 Parsing TeX output . . . . . . . . . . . . . . . . . . . . . . . . . . 48 PATH in Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 PDF mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 Personal customization . . . . . . . . . . . . . . . . . . . . . . . 16 Personal information . . . . . . . . . . . . . . . . . . . . . . . . . 60 Personal macro directory . . . . . . . . . . . . . . . . . . . . . 60 Personal TeX macro directory . . . . . . . . . . . . . . . . 60 pLaTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Poland . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Prefix for labels . . . . . . . . . . . . . . . . . . . . . . . . . . 25, 26 preview-install-styles . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Previewing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 Printing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Private directories . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 Private macro directory . . . . . . . . . . . . . . . . . . . . . . 60 Private style hook directory . . . . . . . . . . . . . . . . . . 60 Private TeX macro directory. . . . . . . . . . . . . . . . . . 60 Problems. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Processes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 pTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55, 57

Q

Quotes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Quotes, fontification of . . . . . . . . . . . . . . . . . . . . . . . 39

R

Redisplay output . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Refilling. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Reformatting . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31, 32 Region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Region file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Reindenting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 Reveal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 Right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Running BibTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Running chktex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Running commands . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Running lacheck . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Running LaTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Running makeindex . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Running TeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

S

Sample style file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Sectioning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 23 Sectioning commands, fontification of . . . . . . . . . 36 Sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 23, 43 Setting the default command . . . . . . . . . . . . . . . . . 44 Setting the header . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Setting the trailer. . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Site customization . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Site information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 Site initialization . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Site macro directory . . . . . . . . . . . . . . . . . . . . . . . . . 59 Site TeX macro directory . . . . . . . . . . . . . . . . . . . . . 59 Slovakia. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Source specials . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 Specifying a font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Starting a previewer . . . . . . . . . . . . . . . . . . . . . . . . . 46 Stopping a process . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 ‘style’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Style file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Style files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Style hook . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Style hooks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Subscript, fontification of . . . . . . . . . . . . . . . . . . . . . 39 Superscript, fontification of . . . . . . . . . . . . . . . . . . . 39 Sweden . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Syntax Highlighting. . . . . . . . . . . . . . . . . . . . . . . . . . 35

T

Tabify . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Table environment . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Tabs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 TeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 TeX parsing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 ‘tex-jp.el’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 ‘tex-site.el’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 tool bar, toolbar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 Trailer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

U

Underfull boxes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 UNICODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Untabify . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 Updating style hooks. . . . . . . . . . . . . . . . . . . . . . . . . 60

V

Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 Verbatim, fontification of . . . . . . . . . . . . . . . . . . . . . 39 Viewing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

W

Warranty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Writing to a printer . . . . . . . . . . . . . . . . . . . . . . . . . . 44

X

‘x-compose.el’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 X-Symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55